### v1

# Adversarial Reasoning for Repair Based on Inferred Program Intent

Abstract.
Automated program repair (APR) has shown promising results, particularly with the use of neural networks. Currently, most APR tools focus on code transformations specified by test suites, rather than reasoning about the program’s intent and the high-level bug specification. Without a proper understanding of program intent, these tools tend to generate patches that overfit incomplete test suites and fail to reflect the developer’s intentions. However, reasoning about program intent is challenging.
In our work, we propose an approach called AdverIntent-Agent, based on critique and adversarial reasoning. Our approach is novel to shift the focus from generating multiple APR patches to inferring multiple potential program intents. Ideally, we aim to infer intents that are, to some extent, adversarial to each other, maximizing the probability that at least one aligns closely with the developer’s original intent. AdverIntent-Agent is a multi-agent approach consisting of three agents: a reasoning agent, a test agent, and a repair agent. First, the reasoning agent generates adversarial program intents along with the corresponding faulty statements. Next, the test agent produces adversarial test cases that align with each inferred intent, constructing oracles that use the same inputs but have different expected outputs. Finally, the repair agent uses dynamic and precise LLM prompts to generate patches that satisfy both the inferred program intent and the generated tests.
AdverIntent-Agent was evaluated on two benchmarks: Defects4J 2.0 and HumanEval-Java. AdverIntent-Agent correctly repaired 77 and 105 bugs in both benchmarks, respectively. Our work suggests a new way for developers to decide whether to accept a patch by providing a complete package of program intent, tests, and code changes. This helps reduce the effort needed to review patches, going beyond just code transformation and test adequacy.
1. Introduction
Automated program repair (APR) aims to reduce the manual and costly processes involved in software maintenance tasks related to bug fixing (Monperrus2015, ; TSE-repair-survey, ). A wide range of approaches have been proposed to modify programs at the source level to bring behavior in line with a given test suite. These techniques include search-based (LeGoues2012GenProg, ; ssFix, ; sofix, ; Yuan2017ARJAAR, ; astor, ; ali-issta19-bytecode, ; sharpFix, ; relifix-shinwei-icse15, ; APR-Fuzzing-ISSTA23, ), semantics-based (Angelixicse16, ; acs, ; JFix, ; concolic-repair-PLDI21, ; patch-transplantation-TOSEM21, ; CrashProgramRepair-ISSTA19, ), and learning-based (codit-tse20, ; SEQUENCER, ; RewardRepair-icse22, ; Tufano-ICSE19, ; modit-ase21, ; CoCoNuT, ; Recoder, ; ye2022selfapr, ; ITER, ) techniques. Not unsurprisingly, the emergence of Large Language Models (LLMs) has motivated growing interest in using program repair techniques, such as code generated by GitHub Copilot (CopilotingCopilots, ), Amazon Code Whisperer (yeticstiren2023evaluating, ), and Replit (fan-automatedicse23, ). Recent large-scale industrial deployments from Google (googlerepair, ) and Meta (sapfix, ) provide promising initial validation that techniques developed in the lab can fruitfully move to industrial practice.
However, existing deployments (and research prototypes) remain limited in the types of bugs they successfully target. Meta’s SapFix, for example, specifically targets certain types of null pointer exceptions. This allows tool designers to have more confidence that narrowly tailored patches are likely to be correct and acceptable. Acceptability in program repair has long been informed by program behavior on test suites, as tests are common in practice, and well-understood. However, overfitting to provided tests (cure-worse-15, ) is a known problem, resulting in patches that cause tests to pass, but do not generalize to the true specification, or developer intent. Researchers have proposed to tackle overfitting in traditional program repair through post-assessment approaches, such as additional heuristics (ICSE18-patchsim, ), probability models (ye2021ods, ; TianASE20, ; Jooyong-Poracle-TOMSE23, ), or test augmentation techniques (drr, ; gao2019crash, ). However, the problem is intrinsic in patch generation, and existing techniques remain limited in the complexity of defects they can empirically correctly tackle.
We argue that such traditional techniques are limited in particular by their focus on transformation types in their design (such as new templates, or new types of synthesis, to use to attempt various types of repair). They do not, by and large, attempt to reason about higher level intent of a given program. An intent in this work is defined as the expected behavior of a function specification that developers aim to achieve. LLM-driven techniques have recently offered an alternative perspective, potentially increasing APR expressiveness and efficiency by both (a) engaging an LLM to reason about a program’s intended behavior, bringing developer intent more explicitly into the process, and (b) iteratively refining initial patch attempts towards more complex patches. This line of LLM-based work enables a self-repair loop to refine previously generated intermediate patches, which is considered state-of-the-art in the field (hidvégi2024cigar, ; ChatRepair-ISSTA24, ; ITER, ).
This idea is promising, but incomplete: iterative, conversational repair approaches like ChatRepair (ChatRepair-ISSTA24, ), ITER (ITER, ), and Cigar (hidvégi2024cigar, ) fundamentally assume that an initial repair attempt proposed by an LLM can be refined to an acceptable solution. However, LLMs often face challenges in reasoning, and the initial bug-fixing intent is not always correct (llm-cannot-self-coorrect, ). In such cases, increasing the number of conversational iterations is unlikely to help (is-selfrepair-a-silver-bullet, ). Indeed, during such iterative processes, these approaches often cause conversational agents to issue refusal responses (wang-acl-2023-know, ; xie2024sorrybenchsystematicallyevaluatinglarge, ).
In this work, we propose an approach, AdverIntent-Agent, consisting of three agents: one for reasoning about program intent, another for generating tests based on that intent, and a third for producing patches that satisfy the inferred program intent. The core idea behind AdverIntent-Agent is to adversarially reason through and construct evidence about the developer intent surrounding the buggy code, and use this reasoning to inform patch construction. At a high level, this means that AdverIntent-Agent uses LLM-based agents to (a) infer the developer intention and localize a defect, (b) generate tests to validate both those inferred intentions, and produced patches, supplementing developer-provided tests, and (c) iterates on generated patches conversationally to construct high-quality repairs. More specifically, AdverIntent-Agent:
Uses adversarial reasoning and testing to construct and identify the most likely correct intents. AdverIntent-Agent uses LLMs to infer multiple possible developer intentions with respect to a buggy function. These program intents are, by construction, intended to be independent and adversarial: that is, ideally, it should be impossible for the function, when repaired, to satisfy more than one. AdverIntent-Agent uses test generation to help demonstrate and validate that the inferred intentions are likely adversarial, assuming the intention is correct by construction and thereby sidestepping the oracle problem (oracleproblem, ). These tests allow for more effective automated reasoning about which intention is correct, and reduces the risk of overfitting to the developer-provided tests. The approach also increases the diversity of the considered patch pool by construction, increasing the likelihood of repair success. In practice, it is difficult to construct completely adversarial intents. Our work is an early exploration that uses test cases to quantify the degree of adversariality and to maximize the distinctiveness of intents within a limited number of attempts.
Dynamically constructs precise prompts to explore alternative bug root causes. AdverIntent-Agent uses adversarial reasoning to explore multiple possible program intentions. However, multiple root causes are still possible, even given a single possible program intent. AdverIntent-Agent uses dynamic precise prompts to first ask LLMs to reason about the top K root causes of bugs based on a given inferred intent. We use those root causes to seed additional prompts to request patches that address each root cause. This approach contrasts with previous uses of generic prompts for patch refinement (ChatRepair-ISSTA24, ), as it explicitly guides the generation of diverse solutions.
We implement AdverIntent-Agent and evaluate it against two benchmarks: Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023knod, ). On the Defects4J 2.0 benchmark, AdverIntent-Agent successfully repairs the highest number of 77 bugs compared to related works (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; repairagent, ; ye2022selfapr, ), and achieving a total of 105 bug repairs in HumanEval-Java, which surpasses all considered related works. These findings highlight the effectiveness of AdverIntent-Agent in addressing the challenges of automated program repair.
We make the following contributions:
  • We devise AdverIntent-Agent, an adversarial reasoning agent for program intents, to guide the generation of diverse and high-quality patches for program repair.
  • AdverIntent-Agent is extensively evaluated on two datasets, including Defects4J v2.0 and HumanEval, achieving state-of-the-art performance.
  • Our work explores an original technique of adversarial program intents, where the degree of adversarial is measured by the number of adversarial test cases that can be generated.
  • Our work fundamentally shifts the interaction with developers regarding patch acceptance by providing a rich suggestion package that includes program intent, tests, and patches.
2. Illustration
Figure 1. An illustrative example AdverIntent-Agent to show the difference between initial and adversarial reasoning and corresponding patches and test case generated.
public int countUpper(String s) {
for(int i=0; i<s.length(); i+=2){
char c = s.charAt(i);
if(c==’A’||c==’e’||c==’I’||c==’o’||c==’u’) {
count += 1;
return count;
((a)) Buggy program from HumanEval-Java
//Test case to expose the bug
public void testCountUpper() {
int result = countUpper(”aBCdEf”);
assertEquals(1, result);
[…]
((b)) Initial failing test case to expose the bug
Initial Program Intent 1: The program is to count every uppercase vowels in ‘s‘. The faulty code is: - for (int i=0; i¡ s.length(); i += 2 { - if (c == ’A’——c == ’e’——c == ’I’——c == ’o’…
Adversarial Program Intent 2: The program is to count uppercase vowels at even indices in string ‘s‘. The faulty code is: - if (c == ’A’——c == ’e’——c == ’I’——c == ’o’…
Adversarial Program Intent 3 The program is to count uppercase and lowercase vowels at even indices in string ‘s‘. - if (c == ’A’——c == ’e’——c == ’I’——c == ’o’…
((c)) Agentreason reasons three program intents
Initial Test Cases 1 : Input: s = ”UNIvERsiTy”, Expected output: 3 Input: s = ”Apple”, Expected output: 1 Input: s = ”bAnaNa”, Expected output: 1 Input: s = ”AeioOU”, Expected output: 3
Adversarial Test Cases 2 : Input: s = ”UNIvERsiTy”, Expected output: 3 Input: s = ”Apple”, Expected output: 1 Input: s = ”bAnaNa”, Expected output: 0 Input: s = ”AeioOU”, Expected output: 2
Adversarial Test Cases 3: Input: s = ”UNIvERsiTy”, Expected output: 3 Input: s = ”Apple”, Expected output: 2 Input: s = ”bAnaNa”, Expected output: 0 Input: s = ”AeioOU”, Expected output: 4
((d)) Agenttest generates three sets of adversarial tests to measure the degree of adversarial in different program intents based on different outputs over all generated tests: the adversarial score between the first and second intent is 50%, the adversarial score between first and third intent is 75%.
Patch for Intent 1 - Overfitting: - for (int i=0; i¡ s.length(); i += 2 { + for (int i=0; i¡ s.length(); i += 1 { - if (c==’A’——c==’e’——c ==’I’——c==’o’——c==’u’) { + if (c==’A’——c==’E’——c ==’I’——c==’O’——c==’U’) {
Patch for Intent 2 - Correct: - if (c==’A’——c==’e’——c==’I’——c==’o’——c==’u’) { + if (c==’A’——c==’E’——c==’I’——c==’O’——c==’U’) {
Patch for Intent 3 - non-plausible: - if (c==’A’——c==’e’——c == ’I’——c==’o’——c==’u’) { + if (c==’A’——c==’E’——c == ’I’——c==’O’——c==’U’ —— c==’a’——c==’e’——c == ’i’——c==’o’—— c==’u’) {
((e)) Agentrepair generates three sets of patches to satisfy program intents and pass adversarial tests.
In this section, we illustrate the workflow of AdverIntent-Agent with an example. Consider 1(e), showing a bug in the CountUpper program from the Humaneval-Java (jiang2023knod, ) dataset. We use this example to demonstrate AdverIntent-Agent’s initial and adversarial approach to formulate program intention as part of a repair process.
Given the buggy program 1(e) and a test failure that exposes the bug 1(e), AdverIntent-Agent works as follows. First, a reasoning agent (Agentreason) starts with an initial prompt to reason about the intent of program, and its faulty statements: Summarize the program intent, and identify faulty locations. This results in the responses in 1(e). This produces an initial inferred intent and potentially buggy lines (first box of 1(e)), which pinpoints the for loop and the if condition are both incorrect, as it only checks every other letter with mixed lowercase and upper case vowels.
Agentreason also constructs an adversarial prompt. This allows AdverIntent-Agent to simultaneously explore a second possible intent. It therefore asks the LLM, “If the previous intent is NOT correct,…”, producing the adversarial intent to count uppercase vowels at every other position (show in the second box of 1(e)). This second prompt identifies the if condition on line 4 as problematic. AdverIntent-Agent continues this process to achieve another adversarial program intent to count both uppercase and lowercase vowels at even indices (box 3). Note that this differentiates our approach from prior agent-based work (repairagent, ; jiang2023knod, ) by considering different possibilities for program intents.
Next, 1(e) illustrates how the testing agent, Agenttest, uses the two adversarial intents to generate new differential test cases that reflect them. The goal of Agenttest is twofold. First, we use tests to measure the degree of adversarial behavior in the inferred program intents. The prompt instructs the LLM to construct tests with the same inputs but expected different outputs for each program intent. A greater number of differentially generated tests based on the same inputs indicates a higher degree of adversarial behavior in the program intents. When the number of tests falls below a predefined threshold, we reject and re-generate the inferred program intent to ensure all intents are adversarial to each other. In our example in 1(e), the adversarial score between the first and second intents is 50%, as two tests produce different results out of four total tests. In contrast, the adversarial score between intent 1 and intent 3 is 75%. Under our settings, all three intents are considered adversarial, as the threshold is set to 33.3% (the metric is introduced in subsection 3.2. Otherwise, the intents will be re-generated until they achieve the expected degree of adversarialness.
The second goal of Agenttest is to generate tests that reflect each intent and specify the correct behavior for patches. Since we assume each individual intent is correct, there is theoretically a solution to an oracle problem. However, practically, ensuring assertion correctness can be challenging. We apply criticism and ranking strategies to alleviate this issue, which are discussed in detail in subsection 3.2 approach. To the best of our knowledge, AdverIntent-Agent is the first technique to generate test cases at a general level during the patch generation process.
The repair agent, Agentrepair, takes the output of both other agents to generate different sets of patches in 1(e). All generated patches are validated by both the original test cases to assess their plausibility, and on the newly generated adversarial tests reflecting inferred program intent. Our observation is that since the patches are adversarial, it is unlikely that all of them will pass the original test suite. Although incomplete, the developer-provided tests are therefore useful for testing adversarial patch plausibility. In our example, the second patch reflecting adversarial intent 2 is correct, while the other two are either overfitting or non-plausible. If only one is a plausible patch, this increases confidence in the inferred intent, and argues for the use of the new tests to supplement the test suite moving forward. If, on the other hand, all patches are plausible, AdverIntent-Agent increases the likelihood that at least one of them will be correct due to adversarial.
3. Approach
Figure 2. Overview of AdverIntent-Agent that shows three agents that are responsible for reasoning -Agentreason, test generation - Agenttest, and patch generation - Agentrepair.
Figure 2 provides an overview of AdverIntent-Agent. AdverIntent-Agent takes as input a program and a set of test cases that identify a bug. It also is able to interact with an execution environment (by, e.g., compiling, executing, or testing the generated code) to collect information as necessitated by the process, or in response to LLM requests.
AdverIntent-Agent then consists of a three-component multi-agent framework:
  • Reasoning Agent (Agentreason): Reasons about multiple program intents at the function level and locates fault statements that violate such intent.
  • Test Agent (Agenttest): Generates test cases based on the inferred program intent, to verify the adversarial degree of generated intent and act as oracles on the repair process.
  • Repair Agent (Agentrepair): Creates patches based on the located faulty statements and program intent.
A core element of AdverIntent-Agent’s approach, as illustrated in 1(e), is the use of adversarial reasoning. The key to achieving reasoning is through criticism prompts to tell LLMs the previous answer is insufficient and explicitly to explore diversity, e.g., The previous answer is not correct, consider alternatives. Each agent thus simultaneously explores solutions to the initial repair task, as well as different adversarial formulations of the task. This approach is novel compared to related work as follows: 1) AdverIntent-Agent is the first work to reason about multiple, partially adversarial program intents, increasing the likelihood that at least one aligns with the correct intent. 2) AdverIntent-Agent is the first work, to our knowledge, to augment test generations during patch generation process, on the opposite of post patch assessments (ICSE18-patchsim, ; Jooyong-Poracle-TOMSE23, ) or specific oracle construction (e.g., crash (gao2019crash, ). 3) AdverIntent-Agent is explicitly guided to explore differences and alternatives, especially compared with prior conversational approaches (ChatRepair-ISSTA24, ; ITER, ), that limited themselves to iteratively refine generated patches on previous answers.
There are two kinds of interactions in AdverIntent-Agent: (1) Multi-Agent interactions and (2) Agent-Environment interactions. In Multi-Agent interactions, Agentreason infers the function intent for Agenttest and identifies faulty statements for Agentrepair. Agenttest generates tests based on the inferred intent to: 1) measure the degree of adversarial behavior between two inferred intents and reject one if the adversarial score falls below a predefined threshold; and 2) use these newly generated adversarial tests to validate the patches produced by Agentrepair. As to Agent-Environment interactions: Agentreason retrieves bug information from the file system to provide additional information requested by the LLMs. Agenttest may retrieve sample test cases from the file system, while Agentrepair requires the environment to validate the generated patches and collect execution feedback.
The rest of this section describes each agent in AdverIntent-Agent and Table 1 shows prompt examples throughout.
3.1. Agentreason
The goal of Agentreason is two-fold: (1) infer program intent to understand expected behavior and (2) locate the buggy statements. As shown in Figure 2, Agentreason infers program intents sequentially, first generating an initial intent and then continuously generating adversarial ones based on previous intents, rather than generating them in parallel, to control each subsequent intent differs from the previous one.
Agentreason takes three inputs: (a) the source code for the presumed-buggy class; (b) the failing test cases; (c) the error messages, which identify the buggy class from the failing test cases where the function is tested but causes test failures. Agentreason works in two steps. First, it narrows down the buggy class to a finer granularity (functions, constructors, and variables). Second, it infers the intended program behavior and further refines the faulty statements based on this finer code, such as the function level. As output, Agentreason provides an initial set of program intents and faulty statements that exist in the current version. Each program intent is expected to be explored adversarially and distinctly.
Table 1. Prompts used in AdverIntent-Agent
Agent Description Prompts
Agentreason Fault Localization (from class level to finer granularity)     P1: Locate the top-3 faulty functions from the given buggy class.
    P2: If the fault does not exist in the function, could you locate the variable definition or other potentially buggy classes?
    P3: What if the previous answers are incorrect? What alternatives are available?
Reason Program Intent     P4: Given the buggy code snippet, can you reason about the program intent and locate the corresponding fault statements?
    P5: What if the previous intents are incorrect? What alternatives are available?
Agenttest Initial Test Generation     P6: Generate N tests based on provided program intent.
Criticism Assertions     P7: Are there any assertions that could be wrong?
Adversarial Test Generation     P8: Modify the previous tests based on this intent to maintain the same inputs but generate different outputs.
Test Prioritization     P9: Rank tests based on confidence of assertion correctness.
Agentrepair Dynamic Precise Prompts     P10: What are the top-3 most likely root causes of this bug-breaking program intent? Answer in X, Y, Z.
    P11: Generate a patch to repair the bug caused by X.
    P12: Generate a patch to repair the bug caused by Y.
    P13: Generate a patch to repair the bug caused by Z.
Patch Refinement     P14: Refine the patches based on the compilation and execution errors.
Adversarial fault localization refined from the class level to statement granularity. Agentreason adversarially explores potential fault localizations by critiquing previous answers and guiding alternative exploration through three adversarial prompts. As shown in the first row of Table 1, Agentreason iteratively narrows the scope of faulty statements from the entire buggy class to finer granularity levels, such as functions (P1). Additionally, Agentreason explicitly guides LLMs to explore statements outside the function level, including constructors and variables, to achieve statement-level fault localization (P2). The LLM is also directed to examine related statements from the initially identified buggy class, as similar issues may exist in other classes, thereby enabling multi-location reasoning (P3). Initially, only a single class file is provided. After adversarial prompting and evaluating the error message, the LLM may request additional information. We then supply the requested files to facilitate cross-class fault localization. This design explicitly and systematically explores a diverse set of faulty statements, complementing previous answers. Note that Agentreason is novel in integrating LLM-based fault localization with program repair and reasoning about program behavior. Prior work either assumed perfect fault localization (e.g., RepairAgent (repairagent, )) and others  (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; CURE-icse21, ; SEQUENCER, ; CoCoNuT, ). Other techniques integrate with spectrum-based fault localization (see, e.g., GenProg (le_goues_genprog_2012, ) and related techniques  (RewardRepair-icse22, ; ye2022selfapr, )). This includes agent-like AutoCodeRover (autocoderover, ), which considers fault localization and repair as two separate tasks and increases configuration effort.
With the identified faulty functions and statements from the first step, Agentreason begins by generating the initial program intent by asking the LLM, ”Can you reason about the program intent and locate the corresponding fault statements?” (P4). Based on the response, Agentreason explicitly guides the LLM to reason about alternative program intents by posing the critical prompt, ”What if the previous intents are incorrect? What alternatives are available?” (P5). These critical prompts can be repeated K times until K adversarial intents are generated. An example is provided earlier in 1(e), where the response summarizes the program intents in natural language along with the corresponding faulty statements that violate those intents. Our insights are twofold. First, Agentreason provides multiple program intents, which are, to some degree, adversarial to each other, thus covering the entire search space of possible intents. This increases the likelihood that at least one of them is correct or close to the original developer’s intent. Second, by assuming each individual intent is correct, this approach partially addresses the oracle problem, enabling test oracles to be constructed as long as they satisfy the inferred program intent.
3.2. Agenttest
Agenttest takes as input program intents descriptions inferred by Agentreason, and produces as output executable test cases to reflect those behaviors. As shown in Figure 2, test generation is a sequential process in which initial tests based on the initial program intent are generated first, followed by tests based on adversarial intents in sequence. Note that each set of test cases is targeted toward validating a particular hypothesis about the function intent.
There are two objectives in Agenttest
First, Agenttest generates tests to measure the adversarial degree of inferred intents. If the adversarial degree of an intent is low, that intent will be removed, and a new adversarial intent will be generated. Second, Agenttest generates additional tests to address the overfitting problem by validating the patches generated by Agentrepair. Specifically, we validate whether the generated patches are overfitting to a specific intent. When one of the intents is close to the developer’s intent, the generated patches are considered correct rather than overfitting. Note that this use of tests generated as an integral part of repair is, to our knowledge, novel. Overfitting patch assessment tends to be considered a post-processing step in patch generation, or left to manual effort (as in previous agent-based approaches (repairagent, ; fix-agent, ; autocoderover, ), which conclude repair attempts upon finding plausible test-passing patches).
Initial test generation and assertion criticism.
In the initial test generation, Agenttest prompts the LLMs to generate N tests based on the initial program intent (P6), establishing a baseline for adversarial tests. However, before modifying these initial tests to create adversarial ones, we use a critique prompt to encourage the LLMs to double-check the correctness of assertions by asking ”Are there any assertions that could be wrong?” (P7). To ensure accurate oracles, we remove any tests for which the LLMs are uncertain about their correctness. Since not all LLM-generated test cases are executable. Agentrepair attempts to iteratively repair generated tests until they compile, up to three times. Test cases that still do not compile after three repair attempts are discarded.
Adversarial test generation
Recall that Agentreason produces multiple adversarial program intents. Similarly, Agenttest aims to generate adversarial tests for these inferred intents. In this context, adversarial tests refer to those with the same input but with expected outputs that vary depending on which intent the program satisfies. Given the initial set of tests, adversarial prompting directs the LLM to “Modify the expected output according to the adversarial function intent” (P8). This allows us to align all test oracles to the same inputs and count the number of differing outputs. Inspired by mutation testing (jia2010analysis-mutation, ), we measure the degree of adversarial conflict in each intent with the score $adversarial\_score=\frac{different\_tests}{all\_tests}$. We define an adversarial threshold as $thres=\frac{100\%}{K}$, where K is the number of inferred intents. For example, in Figure 2, $intents\_1$ and $intents\_2$ produce different outputs in two out of four generated tests, resulting in an $adversarial\_score$ of $50\%$.
This design is based on the need to ensure that the test cases are theoretically mutually exclusive, covering the entire intent space, which is our core idea. We use the number of adversarial tests to measure the quality of the newly generated inferred intents. If any inferred intent falls below the threshold, we repeat the generation process until it exceeds the threshold. Consequently, we increase the likelihood that at least one intent aligns with the correct solution. In this work, we simplify the process by comparing each newly inferred intent with the first generated one.
Test Prioritization.
We conduct test prioritization for each set of test cases based on the confidence of LLM assertions by asking the LLMs to “rank test cases based on confidence of assertion correctness” (P9). From the ranked list, we select only the top percentage of tests and filter out the rest according to the project’s test configuration requirements.
3.3. Agentrepair
Agentrepair takes as input the program intents inferred by Agentreason and the tests generated by Agenttest. It aims to generate patches applicable to the input program that fix the bug by ensuring that both the original and adversarial test cases pass. The key difference between Agentrepair and prior conversation-based repair approaches is that Agentrepair explicitly explores the diversity and differences in the root causes of the bug, whereas prior work either limits themselves in terms of diversity by iterative refining based on a single previous patch (ChatRepair-ISSTA24, ; ITER, ) or increases randomness and cost by using random samples in patch generation through adjusting LLM temperatures (hidvégi2024cigar, ; kong2024contrastrepair, ).
Dynamic Adversarial Prompting and Patch Refinement
Building upon the previously inferred program intent, our goal is to precisely identify the root causes of bugs to guide repair actions. The purpose is to diversify repair actions: even with the same program intent, multiple different factors may cause a bug. By precisely identifying these root causes, we ensure diversity in repair actions and prevent all generated patches from converging on a single repair strategy. To this end, Agentrepair employs a novel dynamic and precise prompt construction (P10) to explicitly guide patch generation. To identify the root causes of a bug, Agentrepair first asks the LLMs to identify the top three most likely issues, denoted as X, Y, and Z. These root causes could include ”Null Checks,” ”Floating Point Precision Issues,” ”Array Index Errors,” ”Type Casting,” and ”Negative Numbers.” These root cause keywords are then passed to the next phase to guide patch generation, ensuring that the process remains explicit and controlled.
Next, based on these identified root causes, Agentrepair sends three separate prompts to guide the LLM in fixing the bug according to each root cause (P11, P12, and P13). Each program’s intent leads to three initial patches. However, not all correct patches can be generated in one attempt. Following prior work on conversational LLM approaches (ITER, ; ChatRepair-ISSTA24, ), Agentrepair refines previous patches by executing them against the test cases. If errors are found, the error messages are collected and provided to the LLM for further patch refinement (P14). The process is configured to allow a maximum of refinement configuration reaches.
3.4. Patch Validation
In patch validation, our approach introduces a novel evaluation of patch correctness using original tests, adversarial tests, and manual analysis.
First, we run all patches against the original test cases. Patches that pass these tests are considered plausible. Each plausible patch, generated from a specific inferred program intent, is expected to pass only the adversarial tests associated with that same intent. Assuming one of the inferred intents is correct, if a patch passes the adversarial tests for that intent, it is considered likely-correct with respect to that intent.
Finally, all likely-correct patches are manually analyzed by comparing them to the human-written patch. If judged to be semantically equivalent, they are labeled as believed-correct patches.
4. Evaluation
In this section, we present the experimental setup to evaluate AdverIntent-Agent. Our experiments address the following research questions:
  • RQ1 (Effectiveness): How well does AdverIntent-Agent perform overall in fixing bugs, including as compared to prior techniques?
  • RQ2 (Adversarial Reasoning): To what extent does adversarial reasoning contribute to AdverIntent-Agent’s effectiveness in fault localization and patch generation?
  • RQ3 (Test Cases): To what extent do the test cases generated by Agenttest contributes to the quality of the generated patches?
  • RQ4 (Cost): What is the cost of AdverIntent-Agent in generating plausible patches?
We first evaluate the overall effectiveness of AdverIntent-Agent in repairing software bugs by comparing its performance with prior related work. Next, we conduct an ablation study to assess the contribution of individual components, such as adversarial reasoning and newly generated test cases, to the overall effectiveness. Finally, we analyze the cost of our approach to provide a comprehensive evaluation of its efficiency.
Table 2. Evaluation Benchmark
Datasets Projects # Bugs
Defects4J 2.0 (Defects4J, ) 17 835
HumanEval-Java (jiang2023knod, ) 164 164
Total 181 999
4.1. Experimental Setup
AdverIntent-Agent’s core functionality is implemented in Python, interfacing with the GPT-4o API endpoint. We use a default temperature of 1 to promote patch diversity. We evaluate all generated patches on a 12-core Intel E5-2690V3 CPU at 3.50 GHz with 32GB of RAM, running on Ubuntu 20.04.3 LTS and utilizing OpenJDK Java 64-Bit Server version 1.8.0_312.
In the experiment configuration, we set the number of inferred intents to $K=3$. We aim to locate the top three faulty functions in the provided buggy class and two alternative answers in the constructions, definitions, and other classes, for a total of five. For test generation, we take 70% of the ranked adversarial tests based on the confidence of assertions and filter out the remaining ones. For patch generation, we identify the top three causes that are inconsistent with the inferred program intent, and we allow each initial patch three rounds of refinement.
4.1.1. Benchmarks
We evaluate AdverIntent-Agent on the widely used benchmarks Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023impact, ). Table 2 summarizes details of the considered benchmarks. Defects4J 2.0 consists of 835 bugs from 17 real-world projects and has been widely used in related work, enabling us to conduct a fair comparison. Additionally, we evaluate HumanEval-Java. In HumanEval-Java, developers converted Python programs and their associated test cases from HumanEval into Java programs and JUnit test cases, as well as some bugs were intentionally introduced into the correct Java programs. HumanEval-Java was released after the data collection used for training GPT-3.5, providing a new benchmark for evaluating the model’s ability to handle Java coding tasks.
4.1.2. Evaluating patch quality
We compare AdverIntent-Agent against the state-of-the-art program repair baselines on open-source projects. We measure AdverIntent-Agent and baseline effectiveness at bug repair according to standard metrics in the APR literature:
  • Plausible: the number of bugs that include at least one patch that makes the original developer-written test cases pass.
  • Correct: the number of bugs that include at least one correct patch, as discussed below.
  • $Top@N$: A metric that evaluates whether at least one of the top-$n$ generated patches is correct. A patch is considered correct if it passes both the original test cases and the automatically generated test cases.
We measure APR effectiveness in terms of the number of both plausible (test-passing) patches and correct patches. Patches that overfit to the tests but fail to generalize to the desired specification are a well-known problem in automatic program repair (cure-worse-15, ; kali-issta15, ). We take multiple approaches to evaluate patch correctness. First, we assess whether the patch is plausible that makes the original test cases pass (kali-issta15, ). If a patch is plausible, we further identify whether it is an exact match to the developer-written ground truth patch with string comparison, ignoring whitespace differences.
For the remaining plausible patches that do not exactly match the ground truth, we employ an LLM to assess their correctness based on whether they pass one of the AI-generated tests. If a patch passes its intended test case, it is considered a likely-correct patch; otherwise, it is considered likely-overfitting. Finally, we manually assess patch correctness following prior work (ChatRepair-ISSTA24, ; acs, ; genpro2009, ; CURE-icse21, ). Two authors independently cross-check the correctness of these patches (CURE-icse21, ; ye2022selfapr, ; jiang2023knod, ; repairagent, ). If both authors agree that a patch is semantically equivalent to the ground truth, it is considered believed-correct. These, along with the exact match patches, form the set of correct patches.
4.1.3. Methodology
Overall effectiveness (RQ1)
We compare AdverIntent-Agent with related work across different repair families, including fine-tuning-based methods: SelfAPR (ye2022selfapr, ), AlphaRepair (alpha-repair-fse22, ), ITER (ITER, ); conversation-based methods: ChatRepair (ChatRepair-ISSTA24, ), Cigar (hidvégi2024cigar, ), contrastRepair (kong2024contrastrepair, ), and a baseline from GPT-4; as well as the agent-based approach: RepairAgent (repairagent, ). We reimplemented experiments for ChatRepair (ChatRepair-ISSTA24, ) based on GPT-4 and also conducted a baseline experiment using GPT-4 without execution feedback. For the remaining approaches, we report quantitative results from the corresponding papers and repositories. Plausible and correct repair results are reported as discussed above. In the experiment assuming perfect fault localization, we follow prior work (repairagent, ; kong2024contrastrepair, ; ChatRepair-ISSTA24, ) by providing the buggy statements and their contextual code to the LLMs. For a more realistic comparison, we provide a complete buggy class tested by the failing test cases, allowing the LLM to achieve finer granularity at both the function and statement levels. Existing related work evaluated on HumanEval-Java is mostly based on LLM approaches. This is because HumanEval-Java was released recently, after the data collection used for training GPT-3.5, meaning that earlier approaches were not able to evaluate it at the time they published their paper.
Adversarial Reasoning (RQ2)
We study the effectiveness of adversarial reasoning by: 1) Measuring the adversarial degree between inferred intents to assess how different and distinct they are. This is based on the number of distinct test oracles generated from the same inputs. The evaluation is performed during the initial intent generation, without regenerating intents if the score falls below the threshold. 2) Comparing the alignment between inferred intents and ground-truth intents extracted from the human-written patch. We first use an LLM to generate the ground-truth intent from the patch, then assess semantic equivalence using the GPT-4o. 3) Evaluating its effectiveness as a fault localization and patch generation technique. For buggy statements spanning multiple locations, identifying any part of the buggy statement is considered correct, as it provides a strong starting point for patch generation. This is justified by prior work (ITER, ), which shows that subsequent buggy statements can be discovered iteratively. For patch generation, we compare the number of successful patches generated from the initial intent and the two adversarial intents to evaluate how adversarial reasoning promotes diversity. All three analyses involve substantial manual effort. To make this feasible, we conduct the study on 300 randomly selected bugs from Defects4J 2.0.
Test Generation (RQ3)
In this research question, we are primarily concerned with the true positives—tests that successfully identify and filter out overfitting patches. RQ3 assesses the quality of test cases generated by Agenttest by evaluating their effectiveness in detecting likely-overfitting patches—that is, the extent to which the newly generated tests, as judged by GPT-4o, induce failures in plausible but incorrect patches. For all patches considered likely-overfitting by the LLM, we finally conduct manual analysis to confirm whether they are truly overfitting.
Cost (RQ4)
In RQ4, we study the cost of AdverIntent-Agent in comparison to related work (ChatRepair-ISSTA24, ; hidvégi2024cigar, ; kong2024contrastrepair, ) that also make extensive use of LLM APIs. Specifically, we examine the token cost (i.e., the average number of tokens used to repair one bug), query cost (i.e., the average number of queries made to repair one bug), and ultimately, the estimated monetary cost.
Table 3. Comparison with State-of-the-art on three evaluation benchmarks: Defect4J 2.0 and HumanEval-Java. A * indicates our reimplementation results.
Approach Perfect FL Realistic FL
plausible correct plausible correct
Defects4J 2.0 (835 bugs)
Fine-tuning-based SelfAPR (ye2022selfapr, ) 130 110 107 67
AlphaRepair (alpha-repair-fse22, ) 109 79 90 50
ITER (ITER, ) - - 119 74
LLM Conversation-based ChatRepair (ChatRepair-ISSTA24, ) - 162 - -
Cigar (hidvégi2024cigar, ) - - 185 69
ContrastRepair (kong2024contrastrepair, ) 201 143 - -
Agent-based RepairAgent (repairagent, ) 186 164 - -
AdverIntent-Agent 180 141 135 77
HumanEval-Java (164 bugs)
LLM Conversation-based GPT-4 * 136 127 124 87
ChatRepair * (ChatRepair-ISSTA24, ) 137 130 126 88
ContrastRepair (kong2024contrastrepair, ) 151 137 - -
Cigar (hidvégi2024cigar, ) - - 152 102
AdverIntent-Agent 155 140 146 105
4.2. RQ1: AdverIntent-Agent Effectiveness
Table 3 compares the effectiveness of AdverIntent-Agent with state-of-the-art in two considered benchmark Defects4J 2.0 and HumanEval-Java. The second column lists the existing APR approaches from three categories: tuning-based, LLM-based and Agent-based approaches. Their effectiveness results respectively based on perfect fault localization and realistic fault localization are given in the third and fourth column groups. The best performance number is highlighted in bold in the table. For example, in the last row of the Defects4J 2.0 benchmark, AdverIntent-Agent generated plausible patches for 135 bugs and correctly repaired 77 of them with realistic fault localization. These numbers increase to 180 and 141, respectively, when assuming perfect fault localization.
Avoid overfitting with stronger patch specification by adversarial tests. AdverIntent-Agent achieves the highest number of correct repairs in both the Defects4J 2.0 benchmark (77 of 835 bugs repaired) and the HumanEval-Java benchmark (105 of 164 bugs repaired) under the realistic fault localization configuration. While AdverIntent-Agent does not produce the highest number of plausible patches, this is due to its stronger validation criteria: patches are assessed against both the original and adversarial test cases, where adversarial tests effectively filter out overfitting patches. In contrast, prior work only validates against original test cases, resulting in a higher number of overfitting patches that cannot be eliminated.
Figure 3 shows patches generated based on the same program intent for the add_elements program in HumanEval-Java, both with and without the use of dynamic, precise prompting to guide patch generation. Figure 3 provides an inferred program intent and identifies the faulty statements. There are two issues in this program. First, the buggy code considers at least two digits, which violates the program description of at most two digits (line 4). Second, the buggy code subtracts instead of adding to the final results. Both of these bugs need to be fixed. Figure 3 displays the initial patch generated by Agentrepair. Although both issues are addressed in the patch, it incorrectly identifies all numbers under 100 as having two digits, failing to account for negative numbers that may exceed two digits. The dynamic prompt, shown in Figure 3, first asks the LLM to identify the root cause of the patch to improve it. One response identifies the root cause as ”negative signs,” and a follow-up precise prompt guides the LLM to generate a patch that accounts for negative signs. This leads to a refined patch being generated. Based on the description, AdverIntent-Agent also generates an additional test case to expose the differing program behavior between the initial patch and the refined patch, where the expected program behavior is ’-4’, and the initial patch’s actual output is ’-16’.
Rich information as output AdverIntent-Agent provides rich information, as shown in Figure 3, including inferred program intent and additional test cases, which offer valuable resources for developers to understand the patches. This is a significant distinction from prior work, which typically provides only patches. To our knowledge, AdverIntent-Agent is the only framework that integrates inferred program intents, tests, and patches. This paradigm shift enhances how developers engage with APR patches — enabling them from understanding the patch and choosing one program intent that most closely aligns with their original intent.
Figure 3. AdverIntent-Agent generates a correct patch based on dynamic explicit root cause guidance
An inferred program intent and located faulty statements: Program Intent: Given an array of integers arr and an integer k, the program should return the sum of elements that have at most two digits, within the first k elements of arr. For this calculation, any negative sign should count as a digit, so negative numbers up to -99 (e.g., -5, -23) qualify as ”two-digit numbers” under this rule. Suspicious Faulty Statements - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug: at least instead of at most - result -= arr.get(i); //bug: Subtraction instead of addition - }
((a)) An inferred program intent and faulty statements
//Bug: -10 will be incorrectly count
[…]
((b)) An initial patch
[…]
Generate a patch to address negative signs.
[…]
Replace Math.abs(num)<100‘, convert the
number to a string and count characters.
((c)) Dynamic precise prompt to guide patch refinement
[…]
//Convert the number to a string
[…]
//Check at most two digits
[…]
//Addition instead of subtraction
[…]
((d)) A refined patch based on precise dynamic prompt in (c)
@Test
public void testNegativeNumbers() {
ArrayList arr = new ArrayList(Arrays.
asList(-1, -12, -3));
int k = 3;
// Expected output: -1 + (-3)
assertEquals(-4, add_elements(arr, k));
((e)) A test case that makes (b) initial patch fail and (d) adversarial patch pass
Answer to RQ1: AdverIntent-Agent repairs the highest number of bugs, with 77 in Defects4J 2.0 and 105 in HumanEval-Java. This success is due to effective adversarial reasoning on program intents and dynamic, root-cause-specific prompt construction. The number of plausible yet overfitting patches is low, as all patches undergo robust validation against both original and adversarial tests. Additionally, AdverIntent-Agent offers comprehensive resources for function reasoning, bug analysis, and test cases, providing a unique package in program repair approaches. This design shifts the paradigm by enabling developers to understand APR patches and select program intents more effectively.
4.3. RQ2: Adversarial Intents Effectiveness
Table 4. Intent Alignment Statistics
Scenario Count Percentage
All three intents are aligned 75 25.0 %
Only the first intent is aligned 38 12.7 %
Only the second intent is aligned 21 7.0 %
Only the third intent is aligned 21 7.0 %
Both the first and second are aligned 56 18.6 %
Both the first and third are aligned 17 5.7 %
Both the second and third are aligned 17 5.7 %
At least one intent is considered aligned (total) 245 81.7 %
None of the three intents is considered aligned 55 18.3 %
Figure 4. Adversarial degree between the inferred intents
First, we quantify the adversarial score between different inferred intents. As shown in Figure 4, the majority of adversarial scores fall within the range of 28.6% to 60.0%. The average adversarial score ranges from 43.8% to 46.5%. This data confirms the adversarial nature between different intents.
Second, we measure the extent to which the generated intents align with the expected program intent. We conduct a quality analysis by manually comparing the inferred intents for 300 bugs against the ground-truth patches. Table 4 presents the analysis, showing how well the inferred intents match the ground truth expected intent. Our analysis reveals that in 18.3% of cases (as shown in the last row), none of the inferred program intents are correct. However, it also highlights the significant benefits of using multiple intents. If only a single inferred intent were used, the best single-intent match (first intent) achieves an accuracy of 62.0% (sum of 25.0%, 12.7%, 18.6%, and 5.7%). In contrast, when leveraging multiple intents, at least one correct match occurs in 81.7% of cases.
Table 5. Adversarial Intent Effectiveness
Projects Fault Localization Patch Generation
Initial Intent + Adversarial Intents Initial Intent + Adversarial Intents
Total 64 75 (13.8%) 36 43 (+19.4%)
Table 5 presents an ablation study on the effectiveness of AdverIntent-Agent with and without adversarial intents in fault localization and patch generation. The last row provides a summary of the study. We can see that with adversarial reasoning, fault localization precision increases by 13.8%, and patch generation improves by 19.4%. This confirms the effectiveness of reasoning adversarial intents for both fault localization and patch generation. We make the following implications.
Fault localization performance improves as adversarial intents encourage exploration at both function and statement levels, extending beyond the provided class. First, adversarial reasoning examines a broader and more varied search space for faulty code, not merely by generating more candidate faulty statements. Our experiment shows that simply increasing the number of candidates does not necessarily improve the final result. Second, adversarial reasoning guides the LLM to locate buggy statements outside functions, such as variable definitions, constructors, and other classes. For example, the bug Math-104, involving a buggy variable definition, was identified only through adversarial prompting and was missed in the initial response, which was limited to the function level. Similarly, the bug Math-22 is fixed as AdverIntent-Agent generates patches across two classes, rather than confining itself to the provided class.
Patch generation performance improves as adversarial intents increase the likelihood of the search space containing a correct patch. We observe that successful patches benefit from prompts that are both adversarial and precise. The number of correctly repaired bugs increased from 36 to 43 by incorporating adversarial intents. Adversarial intents explore a diverse range of expected program behaviors, avoiding repeated flawed attempts and effectively pinpointing root causes with accuracy. This approach is a key difference from prior conversational methods, which focus on limited diversity.
We also observe that adversarial intents greatly benefit from an agent-driven search and decision loop When AdverIntent-Agent creates prompts that encourage the LLM to consider adversarial reasoning, the LLM occasionally requests additional information, such as a new class file. The agent then provides this extra context, enabling exploration of different patch possibilities and helping avoid dead-end refusal responses. Without agent-like interaction, the bug repair loop is more likely to halt prematurely due to adversarial reasoning.
Answer to RQ2: Adversarial intent reasoning significantly boosts fault localization precision by 13.8% and enhances patch generation by 19.4%. Adversarial reasoning expands the search space by encouraging exploration beyond function-level constraints, effectively locating buggy statements and enhancing patch accuracy. Additionally, the agent-driven loop facilitates broader exploration by providing context, helping avoid dead-end responses.
4.4. RQ3: Tests and Overfitting
Table 6. Generated Test cases and Overfitting Alleviation
Test Cases Likely-overfitting Detected
Initial Low-Confidence Rank Threshold Valid
Defects4J 2.0 25050 28.4% 70% 3916 12
HumanEval-Java 4920 23.8% 70% 1666 8
Table 6 shows the number of tests generated by Agenttest and the number of bugs with overfitting patches discarded by adversarial tests, where overfitting is associated with the original tests. The second column shows the initial number of generated tests, while the third column indicates the number of low-confidence assertions, as identified by evaluating the correctness of assertions in generated patches, and we prioritize tests based on the LLM’s assertion confidence. The final valid number of tests and the number of bugs with discarded overfitting patches are given in the sixth and seventh columns, respectively.
We make the following observations from Table 6. First, AdverIntent-Agent discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java, confirming the effectiveness of adversarial testing in alleviating overfitting. Second, test generation suffers from a low compilation rate (see the third column), where the successful compilation rate varies between complex, real-world programs and simpler ones. We found that many common errors stem from missing dependencies that cause non-compilation, suggesting potential integration with existing test improvement work (test-issta23, ; unit-test-issta24, ; Greybox-issta23, ). Third, using criticism prompts to ask LLMs to double-check assertions and identify potentially incorrect ones is effective, helping us filter out 23.8% to 28.4% of low-confidence tests.
Answer to RQ3: AdverIntent-Agent effectively discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java. Test generation faces a low compilation rate, with variation between complex and simpler programs, largely due to missing dependencies. Yet, using criticism prompts to double-check assertions filters out over 20% of low-confidence tests, improving test reliability.
4.5. RQ4: Cost of AdverIntent-Agent
Figure 5 in left shows box plot of the token cost across four approaches: ChatRepair, RepairAgent, Cigar, and AdverIntent-Agent. Each box shows the distribution of token costs, with the minimum, first quartile, median, third quartile, and maximum values marked. The figure in right illustrates the cost in cents associated with the median token cost for each approach based on the estimated cost of model GPT-4o. AdverIntent-Agent uses an average of 248K tokens per bug, resulting in a cost of 7.44 cents. We now explain the reason that AdverIntent-Agent has higher costs than RepairAgent and Cigar but lower costs than ChatRepair.
AdverIntent-Agent shows higher costs than RepairAgent due to the realistic fault localization and test generation processes, while AdverIntent-Agent assumes the perfect fault localization is known. AdverIntent-Agent costs are also higher than Cigar’s, due to the prompts of reasoning program intents and test generation, which are key innovations in our approach. However, considering only patch generation, AdverIntent-Agent is more cost-effective than Cigar, as it requests one patch per program intent and root cause, whereas Cigar samples 50 candidates per iteration.
Nonetheless, AdverIntent-Agent remains less costly than ChatRepair, primarily because ChatRepair’s iterative approach includes all historical data in prompts, leading to high token accumulation as iterations reach hundreds. In contrast, AdverIntent-Agent limits itself to three rounds of test and patch refinements, which significantly reduces the total token cost.
Answer to RQ4: AdverIntent-Agent ’s cost increases due to program intent reasoning and test generation, yet remains manageable because of the fewer number of iterations.
Figure 5. Token and Monetary Cost of Repairing Bugs in Defects4J 2.0
5. Related Work
5.1. Automated Program Repair
Code Transformation. Source code transformation to alter program behavior has been the main research focus of many search-based (sofix, ; genpro2009, ; le_goues_genprog_2012, ; Yuan2017ARJAAR, ; Arja-e-Tosem20, ; ssFix, ; long_staged_2015, ; tbar, ; acs, ; martinez2016astor, ; saha2019harnessing, ), constraint-based (semfix, ; Angelixicse16, ; nopol, ; concolic-repair-PLDI21, ; relifix-shinwei-icse15, ; directfix, ), and learning-based repair (tse24-Search-Generate-Modify, ; codit-tse20, ; CoCoNuT, ; DLFix, ; SEQUENCER, ; CURE-icse21, ; jiang2023knod, ; RewardRepair-icse22, ) techniques for a decade. The emphasis lies in constructing syntactically and semantically correct code transformations, taking into account their expressiveness and effectiveness.
Rich Information Beyond Code. Recently, more work considered information beyond code transformation, by including commit message (modit-ase21, ), compiler and test feedback (ye2022selfapr, ; kong2024contrastrepair, ), domain knowledge (tosem24-domain, ; ye2022selfapr, ) and iterative patch refinement (ITER, ). Chen et al. (chen2024teaching, ) improves LMs code generation accuracy by injecting feedback messages generated by the LM itself based on its code execution results.
Conversational-based Repairs. Recent advancements in Large Language Models (LLMs) have paved the way for innovative approaches to automated program repair (APR). ChatRepair(ChatRepair-ISSTA24, ) pioneered a conversational repair method based on LLMs, iteratively generating patches by collecting bug context information and patch execution results. Cigar(hidvégi2024cigar, ) further optimized the repair process by designing prompts and reboot strategies to generate diverse patches while reducing token consumption. ContrastRepair (kong2024contrastrepair, ) takes an LLM-based iterative approach for APR. Their prompt includes a failing test and a similar passing test, where a passing test is either selected from the existing suite or generated using type-aware mutation. CHATDBG (levin2024chatdbg, ) shifted the focus from fully automated repair to a co-pilot for debugging, integrating LLMs with standard debuggers to enhance their capabilities.
Patch Correctness Assessment. Existing methods for assessing overfitting patches typically classify these patches separately and address them in post-processing. Prior work mainly focuses on static code analysis (ye2021ods, ; TianASE20, ; ASE20Wang, ; anti-pattern, ) and test generation and execution analysis (gao2019crash, ; tian2021failtest, ; ICSE18-patchsim, ; issta17-difftgen, ).
Our work falls into the category of conversational-based repairs. Unlike previous prompting-based repair approaches, ours is the first to address the overfitting patch problem during the patch generation phase rather than as a post-generation step.
5.2. LLM-based Agent for Code
There is a set of parallel LLM-based agents designed for repairing software issues related to our research (yang2024sweagent, ; zhang2024codeagent, ; autocoderover, ). Yang et al. (yang2024sweagent, ) propose SWE-Agent for repairing GitHub issues. AutoCodeRover (autocoderover, ) combines LLMs with advanced code search techniques to address GitHub issues through program modification or patch generation. Zhang et al. (zhang2024codeagent, ) introduce CodeAgent, a novel LLM-based agent framework that employs external tools for repository-level code generation, integrating five programming tools for information retrieval, code symbol navigation, and code testing. Bouzenia et al. (repairagent, ) propose RepairAgent, which treats the LLM as an agent capable of autonomously planning and executing actions to fix bugs by invoking suitable tools. Our work similarly uses a decision-action-planning loop with multiple iterations and dynamic prompts based on command outputs, with the novel component of reasoning about multiple adversarial intents to explore diverse solutions.
5.3. Program Reasoning
Program intent reasoning is a critical area of research within software engineering, encompassing various techniques and methodologies aimed at understanding, analyzing, and improving software programs. One notable approach is symbolic execution (varfix, ; sergey-test-equivalence-tosem18, ), which systematically explores program paths to identify potential errors or vulnerabilities. Formal methods provide mathematical techniques for reasoning about program correctness and behavior (SMT, ). SpecRover by Ruan et al. (SpecRover-icse25, ) is a closely related work that extracts code intents via LLMs. This work shares a similar idea with ours by first extracting program intent and then using it to guide patch generation. However, our approach conceptually generates multiple diverse program intents adversarially, maximizing the likelihood that at least one inferred program is correct, whereas SpecRover infers only the most likely program intent.
In the general setting of machine learning for program reasoning, several works propose Chain-of-Thought (yang2023chainofthought, ), Tree-of-Thought (yao2023treeofthoughts, ) and Graph-of-Thought (Graph-of-Thoughts, ) to logical reasoning code generation tasks by breaking them down into understandable intermediate steps, enabling LLMs to handle each step individually. Unlike prior work, our approach reasons about program intent at the function specification level and incorporates multiple adversarial intents to ensure diverse solutions are explored, which is novel to our knowledge.
6. Threats to Validity
We acknowledge several potential threats to validity: 1) Manual patch validation. The first internal threat arises from the manual validation of believed-correct patches against the reference developer patch. To address this, we carefully examined and discussed each patch, and we make the patches publicly available. 2) Data leakage. Using GPT-4o for AdverIntent-Agent raises data leakage concerns due to its undisclosed training data; however, evaluation on the HumanEval-Java dataset helps strengthen internal validity. 3) Non-deterministic of LLMs. Non-deterministic language model outcomes pose a threat. We mitigate this by evaluating a substantial number of bugs and providing interaction logs for reproducibility and transparency. 4) Limited programming languages. Our experiments are limited to Java projects, which threatens external validity, as results may not generalize to other programming languages or projects. To mitigate this, we include two benchmarks, Defects4J 2.0 and HumanEval-Java, covering 181 diverse projects with various bug types, lengths, and complexities, which strengthens the validity of our findings.
7. Conclusion
We present AdverIntent-Agent, a multi-agent system that advances automated program repair by focusing on inferred program intent through adversarial reasoning. Unlike traditional APR approaches, AdverIntent-Agent generates diverse program intents, adversarial tests, and patches guided by precise root causes to reduce overfitting and improve alignment with developer intent. Evaluations on Defects4J 2.0 and HumanEval-Java show that AdverIntent-Agent repairs more bugs than previous methods, underscoring its potential to enhance APR with intent-driven, developer-oriented solutions. Our work suggests a new paradigm in patch assessment that considers developer acceptance of APR patches, shifting to asking developers to analyze the patches and select the inferred program intent that best aligns with their original intent.
8. Acknowledgment
This work was partially supported by The Wallenberg Foundation and KAW Postdoctoral Scholarship Program - KAW 2022.0368. The computations and data handling were enabled by the supercomputing resource Berzelius provided by National Supercomputer Centre at Linköping University and the KAW foundation: 2024 NAISS GPU Grant Berzelius-2024-131.

### v2

@@
 AdverIntent-Agent correctly repaired 77 and 105 bugs in both benchmarks, respectively.
-Our work suggests a new way for developers to decide whether to accept a patch by providing a complete package of program intent, tests, and code changes.
-This helps reduce the effort needed to review patches, going beyond just code transformation and test adequacy.
+Our work helps reduce the effort required to review patches by enabling developers to assess program intent in natural language, rather than reviewing code patches.
 1. Introduction
-Automated program repair (APR) aims to reduce the manual and costly processes involved in software maintenance tasks related to bug fixing (Monperrus2015, ; TSE-repair-survey, ).
+Automated program repair (APR) aims to reduce the manual and costly processes involved in software maintenance tasks related to bug fixing (Monperrus, 2017; Gazzola et al., 2017).
 A wide range of approaches have been proposed to modify programs at the source level to bring behavior in line with a given test suite.
-These techniques include search-based (LeGoues2012GenProg, ; ssFix, ; sofix, ; Yuan2017ARJAAR, ; astor, ; ali-issta19-bytecode, ; sharpFix, ; relifix-shinwei-icse15, ; APR-Fuzzing-ISSTA23, ), semantics-based (Angelixicse16, ; acs, ; JFix, ; concolic-repair-PLDI21, ; patch-transplantation-TOSEM21, ; CrashProgramRepair-ISSTA19, ), and learning-based (codit-tse20, ; SEQUENCER, ; RewardRepair-icse22, ; Tufano-ICSE19, ; modit-ase21, ; CoCoNuT, ; Recoder, ; ye2022selfapr, ; ITER, ) techniques.
-Not unsurprisingly, the emergence of Large Language Models (LLMs) has motivated growing interest in using program repair techniques, such as code generated by GitHub Copilot (CopilotingCopilots, ), Amazon Code Whisperer (yeticstiren2023evaluating, ), and Replit (fan-automatedicse23, ).
-Recent large-scale industrial deployments from Google (googlerepair, ) and Meta (sapfix, ) provide promising initial validation that techniques developed in the lab can fruitfully move to industrial practice.
+These techniques include search-based (Le Goues et al., 2012a; Xin and Reiss, 2017b; Liu and Zhong, 2018; Yuan and Banzhaf, 2018; Martinez and Monperrus, 2016a; Ghanbari et al., 2019; Xin and Reiss, 2019; Tan and Roychoudhury, 2015; Kim et al., 2023), semantics-based (Mechtaev et al., 2016; Xiong et al., 2017; Le et al., 2017; Shariffdeen et al., 2021a, b; Gao et al., 2019a), and learning-based (Chakraborty et al., 2020; Chen et al., 2019; Ye et al., 2022a; Tufano et al., 2019; Chakraborty and Ray, 2021; Lutellier et al., 2020; Zhu et al., 2021; Ye et al., 2022b; Ye and Monperrus, 2024) techniques.
+Not unsurprisingly, the emergence of Large Language Models (LLMs) has motivated growing interest in using program repair techniques, such as code generated by GitHub Copilot (Wei et al., 2023), Amazon Code Whisperer (Yetiştiren et al., 2023), and Replit (Fan et al., 2023).
+Recent large-scale industrial deployments from Google (Google, 2024) and Meta (Marginean et al., 2019) provide promising initial validation that techniques developed in the lab can fruitfully move to industrial practice.
 However, existing deployments (and research prototypes) remain limited in the types of bugs they successfully target.
@@
 Acceptability in program repair has long been informed by program behavior on test suites, as tests are common in practice, and well-understood.
-However, overfitting to provided tests (cure-worse-15, ) is a known problem, resulting in patches that cause tests to pass, but do not generalize to the true specification, or developer intent.
-Researchers have proposed to tackle overfitting in traditional program repair through post-assessment approaches, such as additional heuristics (ICSE18-patchsim, ), probability models (ye2021ods, ; TianASE20, ; Jooyong-Poracle-TOMSE23, ), or test augmentation techniques (drr, ; gao2019crash, ).
+However, overfitting to provided tests (Smith et al., 2015) is a known problem, resulting in patches that cause tests to pass, but do not generalize to the true specification, or developer intent.
+Researchers have proposed to tackle overfitting in traditional program repair through post-assessment approaches, such as additional heuristics (Xiong et al., 2018), probability models (Ye et al., 2021a; Tian et al., 2020; Ismayilzada et al., 2023), or test augmentation techniques (Ye et al., 2021b; Gao et al., 2019b).
 However, the problem is intrinsic in patch generation, and existing techniques remain limited in the complexity of defects they can empirically correctly tackle.
@@
 LLM-driven techniques have recently offered an alternative perspective, potentially increasing APR expressiveness and efficiency by both (a) engaging an LLM to reason about a program’s intended behavior, bringing developer intent more explicitly into the process, and (b) iteratively refining initial patch attempts towards more complex patches.
-This line of LLM-based work enables a self-repair loop to refine previously generated intermediate patches, which is considered state-of-the-art in the field (hidvégi2024cigar, ; ChatRepair-ISSTA24, ; ITER, ).
-This idea is promising, but incomplete: iterative, conversational repair approaches like ChatRepair (ChatRepair-ISSTA24, ), ITER (ITER, ), and Cigar (hidvégi2024cigar, ) fundamentally assume that an initial repair attempt proposed by an LLM can be refined to an acceptable solution.
-However, LLMs often face challenges in reasoning, and the initial bug-fixing intent is not always correct (llm-cannot-self-coorrect, ).
-In such cases, increasing the number of conversational iterations is unlikely to help (is-selfrepair-a-silver-bullet, ).
-Indeed, during such iterative processes, these approaches often cause conversational agents to issue refusal responses (wang-acl-2023-know, ; xie2024sorrybenchsystematicallyevaluatinglarge, ).
+This line of LLM-based work enables a self-repair loop to refine previously generated intermediate patches, which is considered state-of-the-art in the field (Hidvégi et al., 2024; Xia and Zhang, 2024; Ye and Monperrus, 2024).
+This idea is promising, but incomplete: iterative, conversational repair approaches like ChatRepair (Xia and Zhang, 2024), ITER (Ye and Monperrus, 2024), and Cigar (Hidvégi et al., 2024) fundamentally assume that an initial repair attempt proposed by an LLM can be refined to an acceptable solution.
+However, LLMs often face challenges in reasoning, and the initial bug-fixing intent is not always correct (Huang et al., 2024).
+In such cases, increasing the number of conversational iterations is unlikely to help (Olausson et al., 2024).
+Indeed, during such iterative processes, these approaches often cause conversational agents to issue refusal responses (Wang et al., 2023; Xie et al., 2024).
 In this work, we propose an approach, AdverIntent-Agent, consisting of three agents: one for reasoning about program intent, another for generating tests based on that intent, and a third for producing patches that satisfy the inferred program intent.
@@
 These program intents are, by construction, intended to be independent and adversarial: that is, ideally, it should be impossible for the function, when repaired, to satisfy more than one.
-AdverIntent-Agent uses test generation to help demonstrate and validate that the inferred intentions are likely adversarial, assuming the intention is correct by construction and thereby sidestepping the oracle problem (oracleproblem, ).
+AdverIntent-Agent uses test generation to help demonstrate and validate that the inferred intentions are likely adversarial, assuming the intention is correct by construction and thereby sidestepping the oracle problem (Barr et al., 2015).
 These tests allow for more effective automated reasoning about which intention is correct, and reduces the risk of overfitting to the developer-provided tests.
@@
 Our work is an early exploration that uses test cases to quantify the degree of adversariality and to maximize the distinctiveness of intents within a limited number of attempts.
-Dynamically constructs precise prompts to explore alternative bug root causes.
 AdverIntent-Agent uses adversarial reasoning to explore multiple possible program intentions.
 However, multiple root causes are still possible, even given a single possible program intent.
-AdverIntent-Agent uses dynamic precise prompts to first ask LLMs to reason about the top K root causes of bugs based on a given inferred intent.
+AdverIntent-Agent uses dynamic precise prompts to first ask LLMs to reason about the top-k root causes of bugs based on a given inferred intent.
 We use those root causes to seed additional prompts to request patches that address each root cause.
-This approach contrasts with previous uses of generic prompts for patch refinement (ChatRepair-ISSTA24, ), as it explicitly guides the generation of diverse solutions.
-We implement AdverIntent-Agent and evaluate it against two benchmarks: Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023knod, ).
-On the Defects4J 2.0 benchmark, AdverIntent-Agent successfully repairs the highest number of 77 bugs compared to related works (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; repairagent, ; ye2022selfapr, ), and achieving a total of 105 bug repairs in HumanEval-Java, which surpasses all considered related works.
+This approach contrasts with previous uses of generic prompts for patch refinement (Xia and Zhang, 2024), as it explicitly guides the generation of diverse solutions.
+We implement AdverIntent-Agent and evaluate it against two benchmarks: Defects4J 2.0 (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023b).
+On the Defects4J 2.0 benchmark, AdverIntent-Agent successfully repairs 77 bugs, outperforming related works (Xia and Zhang, 2024, 2022; Bouzenia et al., 2024; Ye et al., 2022b).
+In HumanEval-Java, AdverIntent-Agent achieves 105 bug repairs, also surpassing the considered related works.
 These findings highlight the effectiveness of AdverIntent-Agent in addressing the challenges of automated program repair.
@@
   • Our work explores an original technique of adversarial program intents, where the degree of adversarial is measured by the number of adversarial test cases that can be generated.
-  • Our work fundamentally shifts the interaction with developers regarding patch acceptance by providing a rich suggestion package that includes program intent, tests, and patches.
 2. Illustration
-Figure 1. An illustrative example AdverIntent-Agent to show the difference between initial and adversarial reasoning and corresponding patches and test case generated.
 public int countUpper(String s) {
@@
 Adversarial Program Intent 3 The program is to count uppercase and lowercase vowels at even indices in string ‘s‘. - if (c == ’A’——c == ’e’——c == ’I’——c == ’o’…
-((c)) Agentreason reasons three program intents
+((c)) Agentreason reasons three program intents.
 Initial Test Cases 1 : Input: s = ”UNIvERsiTy”, Expected output: 3 Input: s = ”Apple”, Expected output: 1 Input: s = ”bAnaNa”, Expected output: 1 Input: s = ”AeioOU”, Expected output: 3
@@
 Patch for Intent 3 - non-plausible: - if (c==’A’——c==’e’——c == ’I’——c==’o’——c==’u’) { + if (c==’A’——c==’E’——c == ’I’——c==’O’——c==’U’ —— c==’a’——c==’e’——c == ’i’——c==’o’—— c==’u’) {
-((e)) Agentrepair generates three sets of patches to satisfy program intents and pass adversarial tests.
+((e)) Agentrepair generates three sets of patches to satisfy program intents and pass adversarial tests
+Figure 1. An illustrative example AdverIntent-Agent to show the difference between initial and adversarial reasoning and corresponding patches and test case generated.
 In this section, we illustrate the workflow of AdverIntent-Agent with an example.
-Consider 1(e), showing a bug in the CountUpper program from the Humaneval-Java (jiang2023knod, ) dataset.
+Consider Figure 1, showing a bug in the CountUpper program from the Humaneval-Java (Jiang et al., 2023b) dataset.
 We use this example to demonstrate AdverIntent-Agent’s initial and adversarial approach to formulate program intention as part of a repair process.
-Given the buggy program 1(e) and a test failure that exposes the bug 1(e), AdverIntent-Agent works as follows.
+Given the buggy program and a test failure that exposes the bug, as shown in Figure 1, AdverIntent-Agent works as follows.
 First, a reasoning agent (Agentreason) starts with an initial prompt to reason about the intent of program, and its faulty statements: Summarize the program intent, and identify faulty locations.
-This results in the responses in 1(e).
-This produces an initial inferred intent and potentially buggy lines (first box of 1(e)), which pinpoints the for loop and the if condition are both incorrect, as it only checks every other letter with mixed lowercase and upper case vowels.
+These responses are shown in Figure 1.
+This produces an initial inferred intent and potentially buggy lines (first box of Figure 1), which pinpoints the for loop and the if condition are both incorrect, as it only checks every other letter with mixed lowercase and upper case vowels.
 Agentreason also constructs an adversarial prompt.
 This allows AdverIntent-Agent to simultaneously explore a second possible intent.
-It therefore asks the LLM, “If the previous intent is NOT correct,…”, producing the adversarial intent to count uppercase vowels at every other position (show in the second box of 1(e)).
-This second prompt identifies the if condition on line 4 as problematic.
-AdverIntent-Agent continues this process to achieve another adversarial program intent to count both uppercase and lowercase vowels at even indices (box 3).
-Note that this differentiates our approach from prior agent-based work (repairagent, ; jiang2023knod, ) by considering different possibilities for program intents.
-Next, 1(e) illustrates how the testing agent, Agenttest, uses the two adversarial intents to generate new differential test cases that reflect them.
+It therefore asks the LLM, “If the previous intent is NOT correct,…”, producing the adversarial intent to count uppercase vowels at every other position (shown in the second box of Figure 1).
+This second prompt identifies only the if condition on line 4 as problematic.
+AdverIntent-Agent continues this process to achieve another adversarial program intent to count both uppercase and lowercase vowels at even indices (shown in the third box of Figure 1).
+Note that this differentiates our approach from prior agent-based work (Bouzenia et al., 2024; Jiang et al., 2023b) by considering different possibilities for program intents.
+Next, Figure 1 illustrates how the testing agent, Agenttest, uses the two adversarial intents to generate new differential test cases that reflect them.
 The goal of Agenttest is twofold.
@@
 When the number of tests falls below a predefined threshold, we reject and re-generate the inferred program intent to ensure all intents are adversarial to each other.
-In our example in 1(e), the adversarial score between the first and second intents is 50%, as two tests produce different results out of four total tests.
+In our example in Figure 1, the adversarial score between the first and second intents is 50%, as two tests produce different results out of four total tests.
 In contrast, the adversarial score between intent 1 and intent 3 is 75%.
@@
 To the best of our knowledge, AdverIntent-Agent is the first technique to generate test cases at a general level during the patch generation process.
-The repair agent, Agentrepair, takes the output of both other agents to generate different sets of patches in 1(e).
+The repair agent, Agentrepair, takes the output of both other agents to generate different sets of patches in Figure 1.
 All generated patches are validated by both the original test cases to assess their plausibility, and on the newly generated adversarial tests reflecting inferred program intent.
@@
 Although incomplete, the developer-provided tests are therefore useful for testing adversarial patch plausibility.
-In our example, the second patch reflecting adversarial intent 2 is correct, while the other two are either overfitting or non-plausible.
+In our example, the second patch reflecting adversarial intent (middle box of Figure 1) is correct, while the other two are either overfitting (first box of Figure 1) or non-plausible (third box of Figure 1).
 If only one is a plausible patch, this increases confidence in the inferred intent, and argues for the use of the new tests to supplement the test suite moving forward.
@@
 3. Approach
-Figure 2. Overview of AdverIntent-Agent that shows three agents that are responsible for reasoning -Agentreason, test generation - Agenttest, and patch generation - Agentrepair.
+Figure 2.
+Overview of AdverIntent-Agent that shows three agents that are responsible for reasoning -Agentreason, test generation - Agenttest, and patch generation - Agentrepair.
+AdverIntent-Agent begins with a buggy class and failing tests as inputs, and its outputs include new intents, tests, and candidate patches.
 Figure 2 provides an overview of AdverIntent-Agent.
@@
   • Repair Agent (Agentrepair): Creates patches based on the located faulty statements and program intent.
-A core element of AdverIntent-Agent’s approach, as illustrated in 1(e), is the use of adversarial reasoning.
+A core element of AdverIntent-Agent’s approach, as illustrated in the motivating example, is the use of adversarial reasoning.
 The key to achieving reasoning is through criticism prompts to tell LLMs the previous answer is insufficient and explicitly to explore diversity, e.g., The previous answer is not correct, consider alternatives.
@@
 This approach is novel compared to related work as follows: 1) AdverIntent-Agent is the first work to reason about multiple, partially adversarial program intents, increasing the likelihood that at least one aligns with the correct intent.
-2) AdverIntent-Agent is the first work, to our knowledge, to augment test generations during patch generation process, on the opposite of post patch assessments (ICSE18-patchsim, ; Jooyong-Poracle-TOMSE23, ) or specific oracle construction (e.g., crash (gao2019crash, ).
-3) AdverIntent-Agent is explicitly guided to explore differences and alternatives, especially compared with prior conversational approaches (ChatRepair-ISSTA24, ; ITER, ), that limited themselves to iteratively refine generated patches on previous answers.
+2) AdverIntent-Agent is the first work, to our knowledge, to augment test generations during patch generation process, on the opposite of post patch assessments (Xiong et al., 2018; Ismayilzada et al., 2023) or specific oracle construction (e.g., crash (Gao et al., 2019b).
+3) AdverIntent-Agent is explicitly guided to explore differences and alternatives, especially compared with prior conversational approaches (Xia and Zhang, 2024; Ye and Monperrus, 2024), that limited themselves to iteratively refine generated patches on previous answers.
 There are two kinds of interactions in AdverIntent-Agent: (1) Multi-Agent interactions and (2) Agent-Environment interactions.
@@
 Each program intent is expected to be explored adversarially and distinctly.
-Table 1. Prompts used in AdverIntent-Agent
+Table 1. Prompts used in AdverIntent-Agent.
 Agent Description Prompts
@@
 Note that Agentreason is novel in integrating LLM-based fault localization with program repair and reasoning about program behavior.
-Prior work either assumed perfect fault localization (e.g., RepairAgent (repairagent, )) and others  (ChatRepair-ISSTA24, ; alpha-repair-fse22, ; CURE-icse21, ; SEQUENCER, ; CoCoNuT, ).
-Other techniques integrate with spectrum-based fault localization (see, e.g., GenProg (le_goues_genprog_2012, ) and related techniques  (RewardRepair-icse22, ; ye2022selfapr, )).
-This includes agent-like AutoCodeRover (autocoderover, ), which considers fault localization and repair as two separate tasks and increases configuration effort.
+Prior work either assumed perfect fault localization (e.g., RepairAgent (Bouzenia et al., 2024)) and others  (Xia and Zhang, 2024, 2022; Jiang et al., 2021; Chen et al., 2019; Lutellier et al., 2020).
+Other techniques integrate with spectrum-based fault localization (see, e.g., GenProg (Le Goues et al., 2012b) and related techniques  (Ye et al., 2022a, b)).
+This includes agent-like AutoCodeRover (Zhang et al., 2024b), which considers fault localization and repair as two separate tasks and increases configuration effort.
 With the identified faulty functions and statements from the first step, Agentreason begins by generating the initial program intent by asking the LLM, ”Can you reason about the program intent and locate the corresponding fault statements?” (P4).
@@
 These critical prompts can be repeated K times until K adversarial intents are generated.
-An example is provided earlier in 1(e), where the response summarizes the program intents in natural language along with the corresponding faulty statements that violate those intents.
+An example is provided earlier in Figure 1, where the response summarizes the program intents in natural language along with the corresponding faulty statements that violate those intents.
 Our insights are twofold.
@@
 Note that this use of tests generated as an integral part of repair is, to our knowledge, novel.
-Overfitting patch assessment tends to be considered a post-processing step in patch generation, or left to manual effort (as in previous agent-based approaches (repairagent, ; fix-agent, ; autocoderover, ), which conclude repair attempts upon finding plausible test-passing patches).
+Overfitting patch assessment tends to be considered a post-processing step in patch generation, or left to manual effort (as in previous agent-based approaches (Bouzenia et al., 2024; Lee et al., 2024; Zhang et al., 2024b), which conclude repair attempts upon finding plausible test-passing patches).
 Initial test generation and assertion criticism.
@@
 This allows us to align all test oracles to the same inputs and count the number of differing outputs.
-Inspired by mutation testing (jia2010analysis-mutation, ), we measure the degree of adversarial conflict in each intent with the score $adversarial\_score=\frac{different\_tests}{all\_tests}$.
+Inspired by mutation testing (Jia and Harman, 2010), we measure the degree of adversarial conflict in each intent with $adversarial\_score=\frac{different\_tests}{all\_tests}$.
 We define an adversarial threshold as $thres=\frac{100\%}{K}$, where K is the number of inferred intents.
@@
 It aims to generate patches applicable to the input program that fix the bug by ensuring that both the original and adversarial test cases pass.
-The key difference between Agentrepair and prior conversation-based repair approaches is that Agentrepair explicitly explores the diversity and differences in the root causes of the bug, whereas prior work either limits themselves in terms of diversity by iterative refining based on a single previous patch (ChatRepair-ISSTA24, ; ITER, ) or increases randomness and cost by using random samples in patch generation through adjusting LLM temperatures (hidvégi2024cigar, ; kong2024contrastrepair, ).
+The key difference between Agentrepair and prior conversation-based repair approaches is that Agentrepair explicitly explores the diversity and differences in the root causes of the bug, whereas prior work either limits themselves in terms of diversity by iterative refining based on a single previous patch (Xia and Zhang, 2024; Ye and Monperrus, 2024) or increases randomness and cost by using random samples in patch generation through adjusting LLM temperatures (Hidvégi et al., 2024; Kong et al., 2024).
 Dynamic Adversarial Prompting and Patch Refinement
@@
 However, not all correct patches can be generated in one attempt.
-Following prior work on conversational LLM approaches (ITER, ; ChatRepair-ISSTA24, ), Agentrepair refines previous patches by executing them against the test cases.
+Following prior work on conversational LLM approaches (Ye and Monperrus, 2024; Xia and Zhang, 2024), Agentrepair refines previous patches by executing them against the test cases.
 If errors are found, the error messages are collected and provided to the LLM for further patch refinement (P14).
 The process is configured to allow a maximum of refinement configuration reaches.
-3.4. Patch Validation
-In patch validation, our approach introduces a novel evaluation of patch correctness using original tests, adversarial tests, and manual analysis.
-First, we run all patches against the original test cases.
-Patches that pass these tests are considered plausible.
-Each plausible patch, generated from a specific inferred program intent, is expected to pass only the adversarial tests associated with that same intent.
-Assuming one of the inferred intents is correct, if a patch passes the adversarial tests for that intent, it is considered likely-correct with respect to that intent.
-Finally, all likely-correct patches are manually analyzed by comparing them to the human-written patch. If judged to be semantically equivalent, they are labeled as believed-correct patches.
 4. Evaluation
@@
   • RQ3 (Test Cases): To what extent do the test cases generated by Agenttest contributes to the quality of the generated patches?
-  • RQ4 (Cost): What is the cost of AdverIntent-Agent in generating plausible patches?
+  • RQ4 (Cost): What is the token cost of AdverIntent-Agent in generating plausible patches?
 We first evaluate the overall effectiveness of AdverIntent-Agent in repairing software bugs by comparing its performance with prior related work.
@@
 Finally, we analyze the cost of our approach to provide a comprehensive evaluation of its efficiency.
-Table 2. Evaluation Benchmark
+Table 2. Evaluation benchmarks.
 Datasets Projects # Bugs
-Defects4J 2.0 (Defects4J, ) 17 835
-HumanEval-Java (jiang2023knod, ) 164 164
+Defects4J 2.0 (Just et al., 2014) 17 835
+HumanEval-Java (Jiang et al., 2023b) 164 164
 Total 181 999
@@
 4.1.1. Benchmarks
-We evaluate AdverIntent-Agent on the widely used benchmarks Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023impact, ).
+We evaluate AdverIntent-Agent on the widely used benchmarks Defects4J 2.0 (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a).
 Table 2 summarizes details of the considered benchmarks.
@@
 HumanEval-Java was released after the data collection used for training GPT-3.5, providing a new benchmark for evaluating the model’s ability to handle Java coding tasks.
-4.1.2. Evaluating patch quality
+4.1.2. Patch Quality Evaluation
 We compare AdverIntent-Agent against the state-of-the-art program repair baselines on open-source projects.
@@
 We measure APR effectiveness in terms of the number of both plausible (test-passing) patches and correct patches.
-Patches that overfit to the tests but fail to generalize to the desired specification are a well-known problem in automatic program repair (cure-worse-15, ; kali-issta15, ).
+Patches that overfit to the tests but fail to generalize to the desired specification are a well-known problem in automatic program repair (Smith et al., 2015; Qi et al., 2015).
 We take multiple approaches to evaluate patch correctness.
-First, we assess whether the patch is plausible that makes the original test cases pass (kali-issta15, ).
+First, we assess whether the patch is plausible that makes the original test cases pass (Qi et al., 2015).
 If a patch is plausible, we further identify whether it is an exact match to the developer-written ground truth patch with string comparison, ignoring whitespace differences.
-For the remaining plausible patches that do not exactly match the ground truth, we employ an LLM to assess their correctness based on whether they pass one of the AI-generated tests.
+For the remaining plausible patches that do not exactly match the ground truth, we employ an LLM to assess their correctness based on whether they pass LLM-generated tests for based on one specific program intent.
 If a patch passes its intended test case, it is considered a likely-correct patch; otherwise, it is considered likely-overfitting.
-Finally, we manually assess patch correctness following prior work (ChatRepair-ISSTA24, ; acs, ; genpro2009, ; CURE-icse21, ).
-Two authors independently cross-check the correctness of these patches (CURE-icse21, ; ye2022selfapr, ; jiang2023knod, ; repairagent, ).
+Finally, we manually assess patch correctness following prior work (Xia and Zhang, 2024; Xiong et al., 2017; Weimer et al., 2009; Jiang et al., 2021).
+Two authors independently cross-check the correctness of these patches (Jiang et al., 2021; Ye et al., 2022b; Jiang et al., 2023b; Bouzenia et al., 2024).
 If both authors agree that a patch is semantically equivalent to the ground truth, it is considered believed-correct.
@@
 4.1.3. Methodology
+We discuss our methodology for each research question below.
 Overall effectiveness (RQ1)
-We compare AdverIntent-Agent with related work across different repair families, including fine-tuning-based methods: SelfAPR (ye2022selfapr, ), AlphaRepair (alpha-repair-fse22, ), ITER (ITER, ); conversation-based methods: ChatRepair (ChatRepair-ISSTA24, ), Cigar (hidvégi2024cigar, ), contrastRepair (kong2024contrastrepair, ), and a baseline from GPT-4; as well as the agent-based approach: RepairAgent (repairagent, ).
-We reimplemented experiments for ChatRepair (ChatRepair-ISSTA24, ) based on GPT-4 and also conducted a baseline experiment using GPT-4 without execution feedback.
+We compare AdverIntent-Agent with related work across different repair families, including fine-tuning-based methods: SelfAPR (Ye et al., 2022b), AlphaRepair (Xia and Zhang, 2022), ITER (Ye and Monperrus, 2024); conversation-based methods: ChatRepair (Xia and Zhang, 2024), Cigar (Hidvégi et al., 2024), ContrastRepair (Kong et al., 2024), and a baseline from GPT-4; as well as the agent-based approach: RepairAgent (Bouzenia et al., 2024).
+We reimplemented experiments for ChatRepair (Xia and Zhang, 2024) based on GPT-4 and also conducted a baseline experiment using GPT-4 without execution feedback.
 For the remaining approaches, we report quantitative results from the corresponding papers and repositories.
 Plausible and correct repair results are reported as discussed above.
-In the experiment assuming perfect fault localization, we follow prior work (repairagent, ; kong2024contrastrepair, ; ChatRepair-ISSTA24, ) by providing the buggy statements and their contextual code to the LLMs.
+In the experiment assuming perfect fault localization, we follow prior work (Bouzenia et al., 2024; Kong et al., 2024; Xia and Zhang, 2024) by providing the buggy statements and their contextual code to the LLMs.
 For a more realistic comparison, we provide a complete buggy class tested by the failing test cases, allowing the LLM to achieve finer granularity at both the function and statement levels.
@@
 For buggy statements spanning multiple locations, identifying any part of the buggy statement is considered correct, as it provides a strong starting point for patch generation.
-This is justified by prior work (ITER, ), which shows that subsequent buggy statements can be discovered iteratively.
+This is justified by prior work (Ye and Monperrus, 2024), which shows that subsequent buggy statements can be discovered iteratively.
 For patch generation, we compare the number of successful patches generated from the initial intent and the two adversarial intents to evaluate how adversarial reasoning promotes diversity.
@@
 Test Generation (RQ3)
-In this research question, we are primarily concerned with the true positives—tests that successfully identify and filter out overfitting patches.
-RQ3 assesses the quality of test cases generated by Agenttest by evaluating their effectiveness in detecting likely-overfitting patches—that is, the extent to which the newly generated tests, as judged by GPT-4o, induce failures in plausible but incorrect patches.
-For all patches considered likely-overfitting by the LLM, we finally conduct manual analysis to confirm whether they are truly overfitting.
+In this research question, we explore to what extent newly generated tests can successfully identify and filter out overfitting patches.
+To evaluate this, we adopt a two-step approach that combines LLM-based selection with manual confirmation.
+Specifically, we employ GPT-4o, which takes as input a set of plausible patches and a set of newly generated tests.
+First, we perform a sanity check on the generated tests by assessing their correctness and removing those with low-confidence oracles.
+Next, we rank the tests and select the top 70% from the ranked list.
+These selected tests are then executed against the plausible patches to identify likely overfitting patches, defined as those that fail the tests.
+For all patches flagged as likely overfitting by the LLM, we conduct a final manual analysis to determine the actual number of true positives.
 Cost (RQ4)
-In RQ4, we study the cost of AdverIntent-Agent in comparison to related work (ChatRepair-ISSTA24, ; hidvégi2024cigar, ; kong2024contrastrepair, ) that also make extensive use of LLM APIs.
-Specifically, we examine the token cost (i.e., the average number of tokens used to repair one bug), query cost (i.e., the average number of queries made to repair one bug), and ultimately, the estimated monetary cost.
-Table 3. Comparison with State-of-the-art on three evaluation benchmarks: Defect4J 2.0 and HumanEval-Java. A * indicates our reimplementation results.
+In RQ4, we conduct a detailed analysis of the cost associated with AdverIntent-Agent by comparing it to several related methods, including ChatRepair(Xia and Zhang, 2024), Cigar (Hidvégi et al., 2024), and RepairAgent (Bouzenia et al., 2024), all of which rely heavily on large language model (LLM) APIs for automated program repair tasks.
+Our comparison focuses specifically on the token cost, which refers to the average number of tokens consumed by each approach to repair a bug in the Defect4J benchmark.
+We evaluate the token cost in real-world software projects, thus we do not evaluate the token cost on HumanEval-Java.
+By evaluating token usage in this context, we aim to provide a clear understanding of the computational resources required by each method and to highlight the efficiency and practicality of AdverIntent-Agent in real-world software engineering scenarios.
+Table 3.
+Comparison with state-of-the-art on three evaluation benchmarks: Defect4J 2.0 and HumanEval-Java.
+A ‘*’ indicates our reimplementation results, and a ‘–’ indicates that the result is not available in the literature.
+The best performance is shown in bold.
 Approach Perfect FL Realistic FL
@@
 Defects4J 2.0 (835 bugs)
-Fine-tuning-based SelfAPR (ye2022selfapr, ) 130 110 107 67
-AlphaRepair (alpha-repair-fse22, ) 109 79 90 50
-ITER (ITER, ) - - 119 74
-LLM Conversation-based ChatRepair (ChatRepair-ISSTA24, ) - 162 - -
-Cigar (hidvégi2024cigar, ) - - 185 69
-ContrastRepair (kong2024contrastrepair, ) 201 143 - -
-Agent-based RepairAgent (repairagent, ) 186 164 - -
+Fine-tuning-based SelfAPR (Ye et al., 2022b) 130 110 107 67
+AlphaRepair (Xia and Zhang, 2022) 109 79 90 50
+ITER (Ye and Monperrus, 2024) - - 119 74
+LLM Conversation-based ChatRepair (Xia and Zhang, 2024) - 162 - -
+Cigar (Hidvégi et al., 2024) - - 185 69
+ContrastRepair (Kong et al., 2024) 201 143 - -
+Agent-based RepairAgent (Bouzenia et al., 2024) 186 164 - -
 AdverIntent-Agent 180 141 135 77
@@
 LLM Conversation-based GPT-4 * 136 127 124 87
-ChatRepair * (ChatRepair-ISSTA24, ) 137 130 126 88
-ContrastRepair (kong2024contrastrepair, ) 151 137 - -
-Cigar (hidvégi2024cigar, ) - - 152 102
-AdverIntent-Agent 155 140 146 105
+ChatRepair * (Xia and Zhang, 2024) 137 130 126 88
+ContrastRepair (Kong et al., 2024) 151 137 - -
+Cigar (Hidvégi et al., 2024) - - 152 102
+AdverIntent-Agent 154 140 146 105
 4.2. RQ1: AdverIntent-Agent Effectiveness
 Table 3 compares the effectiveness of AdverIntent-Agent with state-of-the-art in two considered benchmark Defects4J 2.0 and HumanEval-Java.
-The second column lists the existing APR approaches from three categories: tuning-based, LLM-based and Agent-based approaches.
+The second column lists the existing APR approaches from three categories: fine-tuning-based, LLM-based and Agent-based approaches.
 Their effectiveness results respectively based on perfect fault localization and realistic fault localization are given in the third and fourth column groups.
@@
 For example, in the last row of the Defects4J 2.0 benchmark, AdverIntent-Agent generated plausible patches for 135 bugs and correctly repaired 77 of them with realistic fault localization.
-These numbers increase to 180 and 141, respectively, when assuming perfect fault localization.
-Avoid overfitting with stronger patch specification by adversarial tests.
 AdverIntent-Agent achieves the highest number of correct repairs in both the Defects4J 2.0 benchmark (77 of 835 bugs repaired) and the HumanEval-Java benchmark (105 of 164 bugs repaired) under the realistic fault localization configuration.
+For Defects4J 2.0 benchmark, these numbers increase to 180 and 141, respectively, when assuming perfect fault localization is known.
+For HumanEval-Java benchmark, these numbers increase to 154 and 140, respectively, when assuming perfect fault localization is known.
+Our result shows the overall effectiveness of AdverIntent-Agent.
 While AdverIntent-Agent does not produce the highest number of plausible patches, this is due to its stronger validation criteria: patches are assessed against both the original and adversarial test cases, where adversarial tests effectively filter out overfitting patches.
 In contrast, prior work only validates against original test cases, resulting in a higher number of overfitting patches that cannot be eliminated.
-Figure 3 shows patches generated based on the same program intent for the add_elements program in HumanEval-Java, both with and without the use of dynamic, precise prompting to guide patch generation.
-Figure 3 provides an inferred program intent and identifies the faulty statements.
-There are two issues in this program.
-First, the buggy code considers at least two digits, which violates the program description of at most two digits (line 4).
-Second, the buggy code subtracts instead of adding to the final results.
-Both of these bugs need to be fixed.
-Figure 3 displays the initial patch generated by Agentrepair.
-Although both issues are addressed in the patch, it incorrectly identifies all numbers under 100 as having two digits, failing to account for negative numbers that may exceed two digits.
-The dynamic prompt, shown in Figure 3, first asks the LLM to identify the root cause of the patch to improve it.
-One response identifies the root cause as ”negative signs,” and a follow-up precise prompt guides the LLM to generate a patch that accounts for negative signs.
-This leads to a refined patch being generated.
-Based on the description, AdverIntent-Agent also generates an additional test case to expose the differing program behavior between the initial patch and the refined patch, where the expected program behavior is ’-4’, and the initial patch’s actual output is ’-16’.
-Rich information as output AdverIntent-Agent provides rich information, as shown in Figure 3, including inferred program intent and additional test cases, which offer valuable resources for developers to understand the patches.
+We present a case study to demonstrate the effectiveness of AdverIntent-Agent.
+Figure 3 shows patches generated based on the same program intent for the add_elements program in HumanEval-Java.
+Figure 3 provides an initial inferred program intent and identifies the faulty statement in which the given code checks for at least two digits, which violates the inferred intent of at most two digits (see the comment in Figure 3).
+Figure 3 shows the initial patch generated by Agentrepair.
+Although this patch is intended to fix the issue, it still contains two bugs.
+First, the patch incorrectly identifies all numbers under 100 as having two digits, failing to account for negative numbers that may have more than two digits.
+Second, the initial intent and its patch fail to correct the buggy code that subtracts from, rather than adds to, the final result.
+The adversarial intent, shown in Figure 3, critiques the initial intent and patch and infers a new intent that differs from the initial intent.
+This leads to the generation of an adversarial patch, shown in Figure 3.
+Based on the inferred program intent, AdverIntent-Agent also generates an additional test case Figure 3 to reveal the behavioral difference between the initial patch and the adversarial patch: the expected output is ’-4’ (making the adversarial patch pass), while the initial patch’s actual output is ’-16’ (which fails this test case).
+We can see AdverIntent-Agent provides rich information, including inferred program intent and newly generated test cases, which offer valuable resources for developers to understand the patches.
 This is a significant distinction from prior work, which typically provides only patches.
 To our knowledge, AdverIntent-Agent is the only framework that integrates inferred program intents, tests, and patches.
-This paradigm shift enhances how developers engage with APR patches — enabling them from understanding the patch and choosing one program intent that most closely aligns with their original intent.
-Figure 3. AdverIntent-Agent generates a correct patch based on dynamic explicit root cause guidance
+This paradigm shift enhances how developers engage with APR patches, enabling them from understanding the patch and choosing one program intent that most closely aligns with their original intent.
 An inferred program intent and located faulty statements: Program Intent: Given an array of integers arr and an integer k, the program should return the sum of elements that have at most two digits, within the first k elements of arr.
 For this calculation, any negative sign should count as a digit, so negative numbers up to -99 (e.g., -5, -23) qualify as ”two-digit numbers” under this rule.
-Suspicious Faulty Statements - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug: at least instead of at most - result -= arr.get(i); //bug: Subtraction instead of addition - }
+Suspicious Faulty Statements - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug: at least instead of at most - result -= arr.get(i); - }
 ((a)) An inferred program intent and faulty statements
-//Bug: -10 will be incorrectly count
-[…]
-[…]
-[…]
-[…]
+An initial patch: //Bug: -10 will be incorrectly count - if( Integer.toString (arr.get(i)).length() ¿= 2) { + int num = arr.get(i); + if (Math.abs(num) ¡ 100) {
 ((b)) An initial patch
-[…]
-[…]
-[…]
-Generate a patch to address negative signs.
-[…]
-Replace Math.abs(num)<100‘, convert the
-number to a string and count characters.
-((c)) Dynamic precise prompt to guide patch refinement
-[…]
-//Convert the number to a string
-[…]
-[…]
-//Check at most two digits
-[…]
-//Addition instead of subtraction
-[…]
-[…]
-((d)) A refined patch based on precise dynamic prompt in (c)
-@Test
-public void testNegativeNumbers() {
-ArrayList arr = new ArrayList(Arrays.
-asList(-1, -12, -3));
-int k = 3;
-// Expected output: -1 + (-3)
-assertEquals(-4, add_elements(arr, k));
+An adversarial inferred program intent: //bug 1: at least instead of at most - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug 2: Subtraction instead of addition - result -= arr.get(i);
+((c)) An adversarial inferred program intent and faulty statements
+An adversarial patch: + int num = arr.get(i); + String numStr = Integer.toString(num); + int digitCount = numStr.length(); //Check at most two digits + if (digitCount ¡= 2) { + result += num; //Addition instead of subtraction + }
+((d)) An adversarial patch
+A generated test case: @Test public void testAddElements() { ArrayList arr = new ArrayList(Arrays.
+asList(-1, -12, -3)); int k = 3; // Expected output: -1 + (-3) assertEquals(-4, add_elements(arr, k)); }
 ((e)) A test case that makes (b) initial patch fail and (d) adversarial patch pass
-Answer to RQ1: AdverIntent-Agent repairs the highest number of bugs, with 77 in Defects4J 2.0 and 105 in HumanEval-Java.
+Figure 3. A correct patch generated by AdverIntent-Agent.
+Answer to RQ1: AdverIntent-Agent repairs 77 in Defects4J and 105 in HumanEval-Java benchmarks.
 This success is due to effective adversarial reasoning on program intents and dynamic, root-cause-specific prompt construction.
-The number of plausible yet overfitting patches is low, as all patches undergo robust validation against both original and adversarial tests.
 Additionally, AdverIntent-Agent offers comprehensive resources for function reasoning, bug analysis, and test cases, providing a unique package in program repair approaches.
 This design shifts the paradigm by enabling developers to understand APR patches and select program intents more effectively.
-4.3. RQ2: Adversarial Intents Effectiveness
-Table 4. Intent Alignment Statistics
+4.3. RQ2: Adversarial Intent Effectiveness
+Table 4. Program intent alignment statistics.
 Scenario Count Percentage
@@
 None of the three intents is considered aligned 55 18.3 %
-Figure 4. Adversarial degree between the inferred intents
+Figure 4. Adversarial degree between the inferred intents.
+We analyze the effectiveness of adversarial intents based on 300 randomly selected bugs from the Defects4J benchmark.
 First, we quantify the adversarial score between different inferred intents.
@@
 This data confirms the adversarial nature between different intents.
-Second, we measure the extent to which the generated intents align with the expected program intent.
-We conduct a quality analysis by manually comparing the inferred intents for 300 bugs against the ground-truth patches.
-Table 4 presents the analysis, showing how well the inferred intents match the ground truth expected intent.
-Our analysis reveals that in 18.3% of cases (as shown in the last row), none of the inferred program intents are correct.
-However, it also highlights the significant benefits of using multiple intents.
-If only a single inferred intent were used, the best single-intent match (first intent) achieves an accuracy of 62.0% (sum of 25.0%, 12.7%, 18.6%, and 5.7%).
-In contrast, when leveraging multiple intents, at least one correct match occurs in 81.7% of cases.
-Table 5. Adversarial Intent Effectiveness
-Projects Fault Localization Patch Generation
-Initial Intent + Adversarial Intents Initial Intent + Adversarial Intents
-Total 64 75 (13.8%) 36 43 (+19.4%)
+Second, we measure the extent to which the generated intents align with the expected program intent (ground truth intent).
+Table 4 presents the analysis, showing how well the inferred intents match the expected ground truth intent.
+As shown in the two last rows of Table 4, our analysis indicates that 81.7% of cases with at least one inferred intent are aligned with the ground-truth intent, while only 18.3% of cases have no correctly inferred program intents.
+Specifically, it highlights the significant benefits of using multiple intents.
+If only a single inferred intent is used, the best single-intent match (the first intent) achieves an accuracy of 62.0% (the sum of 25.0%, 12.7%, 18.6%, and 5.7%).
+In contrast, when leveraging multiple intents, at least one correct match is found in 81.7% of cases.
+The second and third inferred intents complement the first intent, accounting for the remaining 19.7% of matches.
 Table 5 presents an ablation study on the effectiveness of AdverIntent-Agent with and without adversarial intents in fault localization and patch generation.
@@
 Second, adversarial reasoning guides the LLM to locate buggy statements outside functions, such as variable definitions, constructors, and other classes.
-For example, the bug Math-104, involving a buggy variable definition, was identified only through adversarial prompting and was missed in the initial response, which was limited to the function level.
-Similarly, the bug Math-22 is fixed as AdverIntent-Agent generates patches across two classes, rather than confining itself to the provided class.
+Table 5. Adversarial intent effectiveness.
+Projects Fault Localization Patch Generation
+Initial Intent + Adversarial Intents Initial Intent + Adversarial Intents
+Total 64 75 (13.8%) 36 43 (+19.4%)
 Patch generation performance improves as adversarial intents increase the likelihood of the search space containing a correct patch.
@@
 Adversarial intents explore a diverse range of expected program behaviors, avoiding repeated flawed attempts and effectively pinpointing root causes with accuracy.
-This approach is a key difference from prior conversational methods, which focus on limited diversity.
-We also observe that adversarial intents greatly benefit from an agent-driven search and decision loop When AdverIntent-Agent creates prompts that encourage the LLM to consider adversarial reasoning, the LLM occasionally requests additional information, such as a new class file.
-The agent then provides this extra context, enabling exploration of different patch possibilities and helping avoid dead-end refusal responses.
-Without agent-like interaction, the bug repair loop is more likely to halt prematurely due to adversarial reasoning.
 Answer to RQ2: Adversarial intent reasoning significantly boosts fault localization precision by 13.8% and enhances patch generation by 19.4%.
@@
 4.4. RQ3: Tests and Overfitting
-Table 6. Generated Test cases and Overfitting Alleviation
-Test Cases Likely-overfitting Detected
-Initial Low-Confidence Rank Threshold Valid
-Defects4J 2.0 25050 28.4% 70% 3916 12
-HumanEval-Java 4920 23.8% 70% 1666 8
 Table 6 shows the number of tests generated by Agenttest and the number of bugs with overfitting patches discarded by adversarial tests, where overfitting is associated with the original tests.
 The second column shows the initial number of generated tests, while the third column indicates the number of low-confidence assertions, as identified by evaluating the correctness of assertions in generated patches, and we prioritize tests based on the LLM’s assertion confidence.
-The final valid number of tests and the number of bugs with discarded overfitting patches are given in the sixth and seventh columns, respectively.
+The number of bugs with likely-overfitting patches that are discarded are given in the last column.
 We make the following observations from Table 6.
-First, AdverIntent-Agent discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java, confirming the effectiveness of adversarial testing in alleviating overfitting.
-Second, test generation suffers from a low compilation rate (see the third column), where the successful compilation rate varies between complex, real-world programs and simpler ones.
-We found that many common errors stem from missing dependencies that cause non-compilation, suggesting potential integration with existing test improvement work (test-issta23, ; unit-test-issta24, ; Greybox-issta23, ).
-Third, using criticism prompts to ask LLMs to double-check assertions and identify potentially incorrect ones is effective, helping us filter out 23.8% to 28.4% of low-confidence tests.
-Answer to RQ3: AdverIntent-Agent effectively discards overfitting patches for 12 bugs in Defects4J 2.0 and 8 bugs in HumanEval-Java.
+First, AdverIntent-Agent discards likely-overfitting patches for 12 bugs in Defects4J 2.0 and 7 bugs in HumanEval-Java, confirming the effectiveness of adversarial testing in alleviating overfitting.
+Second, using criticism prompts to ask LLMs to double-check assertions and identify potentially incorrect ones is effective, helping us filter out 23.8% to 28.4% of low-confidence tests.
+Answer to RQ3: AdverIntent-Agent effectively discards overfitting patches for 12 bugs in Defects4J 2.0 and 7 bugs in HumanEval-Java.
 Test generation faces a low compilation rate, with variation between complex and simpler programs, largely due to missing dependencies.
-Yet, using criticism prompts to double-check assertions filters out over 20% of low-confidence tests, improving test reliability.
 4.5. RQ4: Cost of AdverIntent-Agent
-Figure 5 in left shows box plot of the token cost across four approaches: ChatRepair, RepairAgent, Cigar, and AdverIntent-Agent.
-Each box shows the distribution of token costs, with the minimum, first quartile, median, third quartile, and maximum values marked.
-The figure in right illustrates the cost in cents associated with the median token cost for each approach based on the estimated cost of model GPT-4o.
-AdverIntent-Agent uses an average of 248K tokens per bug, resulting in a cost of 7.44 cents.
-We now explain the reason that AdverIntent-Agent has higher costs than RepairAgent and Cigar but lower costs than ChatRepair.
-AdverIntent-Agent shows higher costs than RepairAgent due to the realistic fault localization and test generation processes, while AdverIntent-Agent assumes the perfect fault localization is known.
-AdverIntent-Agent costs are also higher than Cigar’s, due to the prompts of reasoning program intents and test generation, which are key innovations in our approach.
-However, considering only patch generation, AdverIntent-Agent is more cost-effective than Cigar, as it requests one patch per program intent and root cause, whereas Cigar samples 50 candidates per iteration.
-Nonetheless, AdverIntent-Agent remains less costly than ChatRepair, primarily because ChatRepair’s iterative approach includes all historical data in prompts, leading to high token accumulation as iterations reach hundreds.
-In contrast, AdverIntent-Agent limits itself to three rounds of test and patch refinements, which significantly reduces the total token cost.
+Table 7 compares the average token cost across four approaches: Cigar, RepairAgent, ChatRepair, and AdverIntent-Agent.
+On average, AdverIntent-Agent uses 438K tokens per bug.
+Its token usage is higher than both RepairAgent and Cigar but remains lower than ChatRepair.
+The main reason for the increased cost compared to RepairAgent is that AdverIntent-Agent performs realistic fault localization and automated test generation for each bug, with test generation representing the primary contributor to the overall token cost, a step that is not included in RepairAgent’s workflow.
+Compared to Cigar, AdverIntent-Agent ’s higher token cost results from its use of advanced prompts for reasoning about program intent and generating tests, both of which are key innovations in our approach.
+However, if we consider only the patch generation phase, AdverIntent-Agent is actually more cost-effective than Cigar because it requests three patches per program intent and root cause, whereas Cigar samples 50 candidate patches per iteration, leading to a much greater token expenditure for Cigar in this stage.
+Despite these additional costs, AdverIntent-Agent remains more efficient than ChatRepair.
+ChatRepair’s iterative approach incorporates all historical information into each prompt, causing token consumption to increase rapidly as iterations accumulate, sometimes reaching into the hundreds.
+By contrast, AdverIntent-Agent is designed to limit itself to three rounds of test and patch refinements, which manages overall token usage
+AdverIntent-Agent achieves a balance between realism and efficiency, using more tokens for enhanced reasoning and test generation than some baselines, but remaining substantially more cost-effective than highly iterative methods like ChatRepair.
+Table 6. Generated test cases and overfitting detection.
+Test Cases Likely-overfitting
+Initial Low-Confidence Rank Threshold
+Defects4J 2.0 25050 28.4% 70% 12
+HumanEval-Java 4920 23.8% 70% 7
 Answer to RQ4: AdverIntent-Agent ’s cost increases due to program intent reasoning and test generation, yet remains manageable because of the fewer number of iterations.
-Figure 5. Token and Monetary Cost of Repairing Bugs in Defects4J 2.0
 5. Related Work
@@
 Code Transformation.
-Source code transformation to alter program behavior has been the main research focus of many search-based (sofix, ; genpro2009, ; le_goues_genprog_2012, ; Yuan2017ARJAAR, ; Arja-e-Tosem20, ; ssFix, ; long_staged_2015, ; tbar, ; acs, ; martinez2016astor, ; saha2019harnessing, ), constraint-based (semfix, ; Angelixicse16, ; nopol, ; concolic-repair-PLDI21, ; relifix-shinwei-icse15, ; directfix, ), and learning-based repair (tse24-Search-Generate-Modify, ; codit-tse20, ; CoCoNuT, ; DLFix, ; SEQUENCER, ; CURE-icse21, ; jiang2023knod, ; RewardRepair-icse22, ) techniques for a decade.
+Source code transformation to alter program behavior has been the main research focus of many search-based (Liu and Zhong, 2018; Weimer et al., 2009; Le Goues et al., 2012b; Yuan and Banzhaf, 2018, 2020; Xin and Reiss, 2017b; Long and Rinard, 2015; Liu et al., 2019; Xiong et al., 2017; Martinez and Monperrus, 2016b; Saha et al., 2019), constraint-based (Nguyen et al., 2013; Mechtaev et al., 2016; Xuan et al., 2016; Shariffdeen et al., 2021a; Tan and Roychoudhury, 2015; Mechtaev et al., 2015), and learning-based repair (Liu et al., 2024; Chakraborty et al., 2020; Lutellier et al., 2020; Li et al., 2020; Chen et al., 2019; Jiang et al., 2021, 2023b; Ye et al., 2022a) techniques for a decade.
 The emphasis lies in constructing syntactically and semantically correct code transformations, taking into account their expressiveness and effectiveness.
 Rich Information Beyond Code.
-Recently, more work considered information beyond code transformation, by including commit message (modit-ase21, ), compiler and test feedback (ye2022selfapr, ; kong2024contrastrepair, ), domain knowledge (tosem24-domain, ; ye2022selfapr, ) and iterative patch refinement (ITER, ).
+Recently, more work considered information beyond code transformation, by including commit message (Chakraborty and Ray, 2021), compiler and test feedback (Ye et al., 2022b; Kong et al., 2024), domain knowledge (Zirak and Hemmati, 2024; Ye et al., 2022b) and iterative patch refinement (Ye and Monperrus, 2024).
 Chen et al.
-(chen2024teaching, ) improves LMs code generation accuracy by injecting feedback messages generated by the LM itself based on its code execution results.
+(Chen et al., 2024) improves LMs code generation accuracy by injecting feedback messages generated by the LM itself based on its code execution results.
 Conversational-based Repairs.
 Recent advancements in Large Language Models (LLMs) have paved the way for innovative approaches to automated program repair (APR).
-ChatRepair(ChatRepair-ISSTA24, ) pioneered a conversational repair method based on LLMs, iteratively generating patches by collecting bug context information and patch execution results.
-Cigar(hidvégi2024cigar, ) further optimized the repair process by designing prompts and reboot strategies to generate diverse patches while reducing token consumption.
-ContrastRepair (kong2024contrastrepair, ) takes an LLM-based iterative approach for APR.
+ChatRepair(Xia and Zhang, 2024) pioneered a conversational repair method based on LLMs, iteratively generating patches by collecting bug context information and patch execution results.
+Cigar(Hidvégi et al., 2024) further optimized the repair process by designing prompts and reboot strategies to generate diverse patches while reducing token consumption.
+ContrastRepair (Kong et al., 2024) takes an LLM-based iterative approach for APR.
 Their prompt includes a failing test and a similar passing test, where a passing test is either selected from the existing suite or generated using type-aware mutation.
-CHATDBG (levin2024chatdbg, ) shifted the focus from fully automated repair to a co-pilot for debugging, integrating LLMs with standard debuggers to enhance their capabilities.
+CHATDBG (Levin et al., 2024) shifted the focus from fully automated repair to a co-pilot for debugging, integrating LLMs with standard debuggers to enhance their capabilities.
 Patch Correctness Assessment.
 Existing methods for assessing overfitting patches typically classify these patches separately and address them in post-processing.
-Prior work mainly focuses on static code analysis (ye2021ods, ; TianASE20, ; ASE20Wang, ; anti-pattern, ) and test generation and execution analysis (gao2019crash, ; tian2021failtest, ; ICSE18-patchsim, ; issta17-difftgen, ).
+Prior work mainly focuses on static code analysis (Ye et al., 2021a; Tian et al., 2020; Wang et al., 2020; Tan et al., 2016) and test generation and execution analysis (Gao et al., 2019b; Tian et al., 2022; Xiong et al., 2018; Xin and Reiss, 2017a).
 Our work falls into the category of conversational-based repairs.
@@
 5.2. LLM-Based Agent for Code
-There is a set of parallel LLM-based agents designed for repairing software issues related to our research (yang2024sweagent, ; zhang2024codeagent, ; autocoderover, ).
+There is a set of parallel LLM-based agents designed for repairing software issues related to our research (Yang et al., 2024; Zhang et al., 2024a, b).
 Yang et al.
-(yang2024sweagent, ) propose SWE-Agent for repairing GitHub issues.
-AutoCodeRover (autocoderover, ) combines LLMs with advanced code search techniques to address GitHub issues through program modification or patch generation.
+(Yang et al., 2024) propose SWE-Agent for repairing GitHub issues.
+AutoCodeRover (Zhang et al., 2024b) combines LLMs with advanced code search techniques to address GitHub issues through program modification or patch generation.
 Zhang et al.
-(zhang2024codeagent, ) introduce CodeAgent, a novel LLM-based agent framework that employs external tools for repository-level code generation, integrating five programming tools for information retrieval, code symbol navigation, and code testing.
+(Zhang et al., 2024a) introduce CodeAgent, a novel LLM-based agent framework that employs external tools for repository-level code generation, integrating five programming tools for information retrieval, code symbol navigation, and code testing.
 Bouzenia et al.
-(repairagent, ) propose RepairAgent, which treats the LLM as an agent capable of autonomously planning and executing actions to fix bugs by invoking suitable tools.
+(Bouzenia et al., 2024) propose RepairAgent, which treats the LLM as an agent capable of autonomously planning and executing actions to fix bugs by invoking suitable tools.
 Our work similarly uses a decision-action-planning loop with multiple iterations and dynamic prompts based on command outputs, with the novel component of reasoning about multiple adversarial intents to explore diverse solutions.
+Table 7. Token cost analysis of AdverIntent-Agent.
+Cigar RepairAgent ChatRepair AdverIntent-Agent
+Tokens 127K 270K 467K 438K
 5.3. Program Reasoning
 Program intent reasoning is a critical area of research within software engineering, encompassing various techniques and methodologies aimed at understanding, analyzing, and improving software programs.
-One notable approach is symbolic execution (varfix, ; sergey-test-equivalence-tosem18, ), which systematically explores program paths to identify potential errors or vulnerabilities.
-Formal methods provide mathematical techniques for reasoning about program correctness and behavior (SMT, ).
+One notable approach is symbolic execution (Wong et al., 2021; Mechtaev et al., 2018), which systematically explores program paths to identify potential errors or vulnerabilities.
+Formal methods provide mathematical techniques for reasoning about program correctness and behavior (Barrett et al., 2009).
 SpecRover by Ruan et al.
-(SpecRover-icse25, ) is a closely related work that extracts code intents via LLMs.
+(Ruan et al., 2025) is a closely related work that extracts code intents via LLMs.
 This work shares a similar idea with ours by first extracting program intent and then using it to guide patch generation.
 However, our approach conceptually generates multiple diverse program intents adversarially, maximizing the likelihood that at least one inferred program is correct, whereas SpecRover infers only the most likely program intent.
-In the general setting of machine learning for program reasoning, several works propose Chain-of-Thought (yang2023chainofthought, ), Tree-of-Thought (yao2023treeofthoughts, ) and Graph-of-Thought (Graph-of-Thoughts, ) to logical reasoning code generation tasks by breaking them down into understandable intermediate steps, enabling LLMs to handle each step individually.
+In the general setting of machine learning for program reasoning, several works propose Chain-of-Thought (Yang et al., 2023), Tree-of-Thought (Yao et al., 2023) and Graph-of-Thought (Besta et al., 2024) to logical reasoning code generation tasks by breaking them down into understandable intermediate steps, enabling LLMs to handle each step individually.
 Unlike prior work, our approach reasons about program intent at the function specification level and incorporates multiple adversarial intents to ensure diverse solutions are explored, which is novel to our knowledge.
@@
 To mitigate this, we include two benchmarks, Defects4J 2.0 and HumanEval-Java, covering 181 diverse projects with various bug types, lengths, and complexities, which strengthens the validity of our findings.
+5) Hallucinations and mistakes of LLMs.
+We use LLMs for test generation based on specific program intent, but the generated oracles can sometimes be incorrect and difficult to detect.
+LLMs are also used to identify overfitting patches, which may result in further mistakes.
+We mitigate this risk by manually reviewing the results of LLM-based detection.
 7. Conclusion
@@
 Our work suggests a new paradigm in patch assessment that considers developer acceptance of APR patches, shifting to asking developers to analyze the patches and select the inferred program intent that best aligns with their original intent.
-8. Acknowledgment
+Acknowledgements.
 This work was partially supported by The Wallenberg Foundation and KAW Postdoctoral Scholarship Program - KAW 2022.0368.